### v1

# Triangular solutions to a Riemann-Hilbert problem from superelliptic curves

Abstract.
We give fundamental solutions of arbitrarily sized matrix Fuchsian linear systems, in the case where the coefficients $B^{(i)}$ of the systems are matrix solutions of the Schlesinger system that are upper triangular, and whose eigenvalues follow an arithmetic progression of a rational difference. The values on the superdiagonals of the matrices $B^{(i)}$ are given by contour integrals of meromorphic differentials defined on Riemann surfaces obtained by compactification of superelliptic curves.
Keywords: Fuchsian linear systems, Riemann-Hilbert problem, isomonodromic deformations, Schlesinger systems, Riemann surfaces, algebraic curves, superelliptic curves.
11footnotetext: Department of mathematics, University of Sherbrooke, Sherbrooke, Quebec, Canada. E-mail: Mouhammed.Anwar.Al.Ghabra@usherbrooke.ca, Vasilisa.Shramchenko@Usherbrooke.ca22footnotetext: Department of mathematics and statistics, University of Montreal, Montreal Quebec, Canada. E-mail: Benjamin.Piche@umontreal.ca
1.  Introduction
Riemann-Hilbert problems are important in mathematics from the theoretical point of view, see the review [17]. At the same time, they play an important role in various applied mathematics contextes, see for example [3, 19, 22, 24] and references therein. In this paper, a particular Riemann-Hilbert problem for a matrix function with Fuchsian singularities is solved through the differential equation satisfied by this matrix function. Namely, we consider a matrix Fuchsian linear problem
(1) $\frac{\mathrm{d}\Phi}{\mathrm{d}z}=\sum_{i=1}^{N}\frac{B^{(i)}}{z-a_{i}}\Phi$
for a $p\times p$ matrix $\Phi(z)$ with $z\in\mathbb{C}P^{1}.$ Here the matrices $B^{(i)}$ are independent of $z$ but may depend on $a_{1},\dots,a_{N}$. Generically, solutions to this equation have singularities at $z=a_{i}$ and $z=\infty$ (unless $\sum_{j=1}^{N}{B^{(i)}}=0$) and, under analytical continuation along an element of $\pi_{1}(\mathbb{C}\setminus\{a_{1},\dots,a_{n}\},z_{0})$, a solution $\Phi(z_{0})$ transforms to another solution related to $\Phi(z_{0})$ by the right multiplication with a constant matrix, called the monodromy matrix corresponding to the chosen loop. This matrix depends only on the homotopy class of the loop. One thus obtains a monodromy representation of the fundamental group of the complex plane punctured at $a_{i}$, that is a homomorphism $\rho:\pi_{1}(\mathbb{C}\setminus\{a_{1},\dots,a_{N}\},z_{0})\to GL(p,\mathbb{C}),$ see [2] The monodromy representation can also be seen as the holonomy representation of the flat meromorphic connection on the trivial rank $p$ vector bundle over $\mathbb{C}P^{1}\setminus\{a_{1},\dots,a_{N},\infty\}$ whose connection form is $\sum_{i=1}^{N}\frac{B^{(i)}}{z-a_{i}}dz.$
Once a fundamental solution to (1) is found, one may obtain its monodromy matrices. The inverse problem, that of reconstructing a function $\Phi(z)$ from its monodromies, is referred to as a Riemann-Hilbert or inverse monodromy problem. Alternatively, it may be reformulated as a problem of finding two matrix functions, each of which is holomorphic either inside or outside of a closed contour passing trough the points $a_{1},\dots,a_{N}$ in the complex plane, such that the two functions are related by the multiplication with a “jump” matrix on the contour. The monodromy matrices associated with the loops around $a_{i}$ may be combined to obtain this jump matrix [4, 17]. The term Riemann-Hilbert problem refers also to the 21st problem from Hilbert’s list, that of reconstructing a Fuchsian equation (1) from a collection of monodromy matrices associated with a set of points $a_{1},\dots,a_{N}\in\mathbb{C}$ [4, 2].
Closely related to the Riemann-Hilbert problems is the search for isomonodromic deformations of Fuchsian systems, that is the question of finding the coefficients $B^{(i)}$ in (1) for which the monodromy matrices of solutions are independent of the small variations of positions of the Fuchsian singularities, that is variations of the set $\{a_{1},\dots,a_{N}\}$. One way of finding such matrices $B^{(i)}$ is by solving the Schlesinger system [25, 5], a system of nonlinear differential equations for $B^{(i)}$ with respect to the variables $a_{1},\dots,a_{N}$ giving a sufficient condition for the resulting family of Fuchsian system to be isomonodromic, that is to share the same monodromies for $a=(a_{1},\dots,a_{N})$ in some set $D\subset\mathbb{C}^{N}$ not intersecting the diagonals of $\mathbb{C}^{N}$.
In the case $p=2$ and $N=3$, setting by a Möbius transformation $\{a_{1},a_{2},a_{3}\}=\{0,1,x\}$, the Schlesinger system reduces to a Painlevé-VI equation with parameters related to the eigenvalues of the matrices $B^{(i)}$, see for example [15].
Solutions to the Schlesinger system are constructed in [6, 20, 8, 9, 10] for $p=2$, in [13, 14] for $p=2$ and $p=3$ and [12, 21, 7] for an arbitrary $p$. Most of these solutions are constructed in terms of functions defined on Riemann surfaces and over families of Riemann surfaces. The surfaces are obtained by considering coverings of $\mathbb{C}P^{1}$ ramified over the set of Fuchsian singularities of the associated system (1), that is over the set $\{a_{1},\dots,a_{N}\}$ and possibly the point at infinity. Allowing the branching set to vary, one considers a family of Riemann surfaces naturally associated with system (1) and thus with the Schlesinger system.
Solving the Fuchsian system (1) having coefficients $B^{(i)}$ obtained as solutions to the Schlesinger system in the algebro-geometric form in the above references is typically quite difficult. One exception being [21] where a Riemann-Hilbert problem with regular singularities was solved for quasi-permutation monodromy matrices (independent of positions of regular singularities), after which the corresponding solution to the Schlesinger system was obtained.
The goal of our paper is to solve system (1) for the coefficients $B^{(i)}$ obtained in [10, 7]. The solutions of the Schlesinger system constructed in [10, 7] are upper triangular and have a relatively simple form due to an additional assumption that the eigenvalues of each matrix $B^{(i)}$ form an arithmetic progression with a rational difference, the difference being the same for all $i=1,\dots,N.$ The eigenvalues of $B^{(i)}$ are called exponents of the associated Fuchsian system. Triangular solutions to Schlesinger systems are also investigated in [11] and [13, 14].
The paper is organized as follows. In Section 2 we recall the construction of matrices $B^{(i)}$ from [7] and the relevant algebraic curves. In Section 3 we present our main result giving solutions to the isomonodromic Fuchsian system with coefficients from [7]. In Section 4 we discuss the monodromy of our solutions and give examples. In Section 5, we consider two particular cases of integration contours leading to rational solutions of the Fuchsian system.
2.  Schlesinger system and superelliptic curves
Let $B^{(1)},B^{(2)},\ldots,B^{(N)}$ be $p\times p$ matrices depending on some complex parameters $a_{1},\ldots,a_{N}$ defined on a ball $D\in\mathbb{C}^{N}\setminus\{(a_{1},\ldots,a_{N})\,\,|\,\,a_{i}=a_{j}\,\text{for some }\,i\neq j\}$. The Schlesinger system for matrices $B^{(j)}$ has the form
(2) $\mathrm{d}B^{(i)}=-\sum_{j=1,j\neq i}^{N}\frac{[B^{(i)},B^{(j)}]}{a_{i}-a_{j}}\mathrm{d}(a_{i}-a_{j}),$
with $i=1,\ldots,N.$
As a partial differential equations system, the Schlesinger system is given by
(3) $\begin{dcases*}\frac{\partial B^{(i)}}{\partial a_{j}}=\frac{[B^{(i)},B^{(j)}]}{a_{i}-a_{j}},\quad i\neq j;\\ \frac{\partial B^{(i)}}{\partial a_{i}}=-\sum_{j=1,j\neq i}^{N}\frac{[B^{(i)},B^{(j)}]}{a_{i}-a_{j}}.\end{dcases*}$
The Schlesinger system was derived [25] as a condition of isomonodromy for the Fuchsian system (1) for a matrix function $\Phi(z)\in M_{p}(\mathbb{C})$. In other words, if the coefficients of system (1), the matrices $B^{(j)}\in M_{p}(\mathbb{C})$, satisfy system (3) as functions of $a_{1},\dots,a_{N}$, then the monodromy matrices of solutions $\Phi$ of (1) on closed loops around $a_{j}$ in the $z$-plane do not change under small perturbations of positions of $a_{1},\dots,a_{N}$.
In [7], some upper triangular matrix solutions to the Schlesinger system (3) were found in terms of contour integrals defined on the compact Riemann surfaces associated with superelliptic curves. In this paper, we give the corresponding upper triangular solutions of the Fuchsian system (1). The superelliptic curves from [7] are given by the equation
(4) $\Gamma_{a}=\{(\zeta,w)\in\mathbb{C}^{2}\,|\,w^{m}=\prod_{i=1}^{N}(\zeta-a_{i})\}$
[…]
$\pi:X_{a}\,\backslash\,\pi^{-1}(\{\infty\})\rightarrow\hat{\Gamma}_{a}\,\backslash\,\{\infty\}$
is biholomorphic. Given that the algebraic curves (4) are non-singular, the map $\pi$ allows us to identify the points of the Riemann surface $X_{a}$ with those of the curve $\Gamma_{a}$ as follows. We say that $P=(\zeta,w)$ is a point of the surface $X_{a}$ if $P=\pi^{-1}([\zeta:w:1]).$ For the points at infinity on $X_{a}$ we use the notation $\{\infty_{1},\dots,\infty_{s}\}:=\pi^{-1}(\{\infty\})$. Note that the number of points at infinity is the greatest common divisor of $N$ and $m$, that is $s={\rm gcd}(m,N)$, and the genus of $X_{a}$ is $g(X_{a})=\frac{1}{2}\bigl{(}(m-1)(N-1)-s+1\bigr{)}$ [7].
The projection to the first coordinate $\zeta$ is a holomorphic function on the algebraic curve $\zeta:\Gamma_{a}\to\mathbb{C}$. It extends naturally to a meromorphic function on the Riemann surface $\zeta:X_{a}\to\mathbb{C}P^{1}$ by setting $\zeta(\zeta,w)=\zeta$ and $\zeta(\infty_{j})=\infty$ with a slight abuse of notation. This way we can identify the Riemann surface $X_{a}$ with the surface of the ramified covering $(X_{a},\zeta)$ of degree $m$ ramified over the points $a_{1},\dots,a_{N},\infty$ called the branch points of the covering. We denote the corresponding ramification points by $P_{a_{i}}=(a_{i},0)$ for $i=1,\dots,N$ and by $\infty_{j}$ as above for $j=1,\dots,s.$
The structure of the ramified covering $\zeta:X_{a}\to\mathbb{C}P^{1}$ induces the standard local coordinates on the surface $X_{a}$ as follows [7]:
$\xi_{a_{k}}(P)=(\zeta(P)-a_{k})^{\frac{1}{m}},$ $\mbox{if}\quad P\sim P_{a_{k}},\quad\mbox{with}\quad k=1,\dots,N\,,$
(5) $\xi_{\infty_{j}}(P)=(\zeta(P))^{-\frac{1}{m_{1}}},$ $\mbox{if}\quad P\sim\infty_{j},\quad\mbox{ where }m=sm_{1}.$
$\xi_{Q}(P)=\zeta(P)-\zeta(Q),$ $\mbox{if}\quad P\sim Q\quad\mbox{and $Q$ is a regular point.}$.
We thus consider the family of compact superelliptic Riemann surfaces $X_{a}$, $a\in D,$ associated with the family of algebraic curves (4). Functions and differentials on the surfaces $X_{a}$ may also be seen as functions over the family of surfaces, that is functions of $a=(a_{1},\dots,a_{N}).$
The solutions from [7] are constructed using the following meromorphic differentials on $X_{a}$ defined for any non-zero integer $n$ coprime to $m$ of (4):
(6) $\Omega_{i}^{j}(a):=\frac{w^{jn}\mathrm{d}\zeta}{\zeta-a_{i}},\quad\text{with}\quad j=1,\ldots,p-1,\quad\text{and}\quad i=1,\dots,N.$
Recall that $\zeta$ is a meromorphic function on $X_{a}$ discussed above. In the same way, $w$ is a meromorphic function on $X_{a}$ extended from $w:\Gamma_{a}\to\mathbb{C}$ by setting $w(\infty_{k})=\infty$ for $k=1,\dots,s.$
Theorem 1.
[7] Let $X_{a}$ with $a=(a_{1},\dots,a_{N})\in D$ be a family of compact Riemann surfaces defined above corresponding to the family of algebraic curves (4). Let $m$ and $N$ be natural integers from (4) and let $n\in\mathbb{Z}$ be coprime to $m$. Let $\Omega_{i}^{j}(a)$ be the meromorphic differential defined by (6) on $X_{a}$. Then a set of upper triangular $p\times p$ matrices $B^{(i)}$, $i=1,...,N$ defined by the following three conditions provides a solution to the Schlesinger system (3):
  • $B^{(i)}_{kl}=0$ if $k>l;$
  • The eigenvalues of each matrix $B^{(i)}$, $i=1,...,N$, satisfy
    (7) $B^{(i)}_{kk}-B^{(i)}_{k+1,k+1}=\frac{n}{m},\quad\text{with }k=1,\ldots,p-1;$
  • The entries on the superdiagonals are given by
    (8) $B^{(i)}_{kl}=\oint_{\gamma_{l-k}}\Omega_{i}^{l-k}(a)=\oint_{\gamma_{l-k}}\frac{w^{(l-k)n}\mathrm{d}\zeta}{\zeta-a_{i}},\quad l>k,$
    where $\gamma_{1},\ldots,\gamma_{p-1}$ are linear combinations with complex coefficients of closed contours on the Riemann surface $X_{a}$ not passing through a pole of the corresponding $\Omega_{i}^{j}(a)$ and independent of small variations of $a\in D$. More precisely, $\gamma_{1},\ldots,\gamma_{p-1}$ are elements of
    1. (a)
      $H_{1}(X_{a},\mathbb{C})$ if $m$, $N$ are coprime and $n>0$,
    2. (b)
      $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s}\},\mathbb{C})$ if $m$, $N$ are not coprime, $s={\rm gcd}(m,N)$, and $n>0$.
    3. (c)
      $H_{1}(X_{a}\setminus\{P_{a_{1}},\ldots,P_{a_{N}}\},{\mathbb{C}})$ if $n<0\,.$
Note that for some contours, the integrals in (8) vanish giving possibly trivial solutions to the Schlesinger system. If $m=1$ and $n>0$ then all the solutions are trivial as well, that is constant diagonal matrices. If the contours are chosen to encircle poles of the respective $\Omega_{i}^{j}(a)$, solutions to the Schlesinger system become rational functions of $a_{1},\dots,a_{N}$, see [7] and Section 5 below.
Solutions given by Theorem 1 are such that $B^{(\infty)}:=\sum_{i=1}^{N}B^{(i)}$ are non-zero constant diagonal matrices, see [7]. This implies that the connection form $\sum_{i=1}^{N}\frac{B^{(i)}}{z-a_{i}}dz$ has a simple pole at $z=\infty$ with residue $B^{(\infty)}.$
3.  Solutions of upper triangular Fuchsian systems
In order to simplify the presentation, let us introduce the following notation.
Definition 2.
Let $n\neq 0$ and $m>0$ be coprime integers, $r=1,\ldots,p-1$, where $p$ is an arbitrary natural number. Let $z\in\mathbb{C}$ be fixed, and let $X_{a}$ be a Riemann surface associated with the superelliptic curve ${\Gamma}_{a}$ (4). Let $\zeta$ and $w=\prod_{i=1}^{N}(\zeta-a_{i})^{1/m}$ be the meromorphic functions on $X_{a}$ as in Section 2, and let $\gamma_{r}$ be a closed contour on $X_{a}$ as in Theorem 1. Assume in addition that $\gamma_{r}$ does not pass by the points $P$ of the surface for which $\zeta(P)=z.$ We then define
(9) $\kappa_{r}:=-\frac{m}{rn}\oint_{\gamma_{r}}\frac{w^{rn}}{\zeta-z}\mathrm{d}\zeta.$
The derivative of $\kappa_{r}$ with respect to $z\in\mathbb{C}$, denoted by $\kappa^{\prime}_{r}$, is obtained in a straightforward way:
(10) $\kappa^{\prime}_{r}:=\frac{\mathrm{d}}{\mathrm{d}z}\kappa_{r}=-\frac{m}{rn}\oint_{\gamma_{r}}\frac{w^{rn}}{(\zeta-z)^{2}}\mathrm{d}\zeta.$
In order to simplify further calculations, let us prove the following lemma.
Lemma 3.
Let $j$ be a strictly positive integer, $\gamma_{j}$ a closed contour on the Riemann surface $X_{a}$, and $\zeta,$ $w=\prod_{i=1}^{N}(\zeta-a_{i})^{1/m}$ meromorphic functions on $X_{a}$ as in Section 2. For $\kappa_{j}$ given by Definition 2 and for its derivative $\kappa^{\prime}_{j}$ (10) with respect to the complex parameter $z$, we have
$\frac{-jn}{m}\kappa_{j}\sum_{i=1}^{N}\frac{1}{z-a_{i}}+\kappa^{\prime}_{j}=\sum_{i=1}^{N}\frac{1}{z-a_{i}}\oint_{\gamma_{j}}\frac{w^{jn}}{\zeta-a_{i}}\mathrm{d}\zeta.$
Proof.   By Definition 2 and (10), we can write
$\frac{-jn}{m}\kappa_{j}\sum_{i=1}^{N}\frac{1}{z-a_{i}}+\kappa^{\prime}_{j}$ $=\sum_{i=1}^{N}\left[\oint_{\gamma_{j}}\frac{w^{jn}\mathrm{d}\zeta}{\zeta-z}\frac{1}{z-a_{i}}\right]-\frac{m}{jn}\oint_{\gamma_{j}}\frac{w^{jn}}{(\zeta-z)^{2}}\mathrm{d}\zeta$
(11) $=\sum_{i=1}^{N}\oint_{\gamma_{j}}\left(\frac{w^{jn}}{\zeta-z}+\frac{w^{jn}}{z-a_{i}}\right)\frac{\mathrm{d}\zeta}{\zeta-a_{i}}-\frac{m}{jn}\oint_{\gamma_{j}}\frac{w^{jn}}{(\zeta-z)^{2}}\mathrm{d}\zeta.$
Due to the vanishing of the integral of an exact differential over a closed contour, for every $z$ we have
$\oint_{\gamma_{j}}\mathrm{d}\frac{w^{jn}}{\zeta-z}=\oint_{\gamma_{j}}\mathrm{d}\frac{\prod_{i=1}^{N}(\zeta-a_{i})^{jn/m}}{\zeta-z}=\oint_{\gamma_{j}}\left(\frac{jn}{m}\frac{w^{jn}}{\zeta-z}\sum_{i=1}^{N}\frac{1}{\zeta-a_{i}}-\frac{w^{jn}}{(\zeta-z)^{2}}\right)\mathrm{d}\zeta=0,$
which together with (11) concludes the proof. $\Box$
The next theorem states the main result of the present paper.
Theorem 4.
Let $p,\;n,\;m,\;N$ and $X_{a},a\in D$ be as in Theorem 1. Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ and let $\gamma_{j}$ be a closed contour on $X_{a}$ from Theorem 1 not passing through any point $P\in X_{a}$ for which $\zeta(P)=z$. More precisely, let $\gamma_{j}$ from Theorem 1 be seen as an element of
  1. (a)
    $H_{1}(X_{a}\setminus\{P\in X_{a}\,|\,\zeta(P)=z\},\mathbb{C})$ if $m$, $N$ are coprime and $n>0$,
  2. (a)
    $H_{1}(X_{a}\setminus(\{\infty_{1},\ldots,\infty_{s}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z\}),\mathbb{C})$ if $n>0$ and $s={\rm gcd}(m,N)>1$,
  3. (b)
    $H_{1}(X_{a}\setminus(\{P_{a_{1}},\ldots,P_{a_{N}}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z\}),{\mathbb{C}})$ if $n<0$,
independent of small variations of $a\in D$. For such a choice of the contour $\gamma_{j}$, let $\kappa_{j}$ be a function of $z$ given by Definition 2. Let $B^{(i)},\;i=1,\dots,N,$ be upper triangular $p\times p$ matrices given by Theorem 1 with eigenvalues
$\beta_{k}^{(i)}=B^{(i)}_{kk}=\theta_{i}-\frac{(k-1)n}{m},\quad k\in\{1,\ldots,p\},$
where $\theta_{i}\in\mathbb{R}$. Then, a fundamental upper triangular matrix solution $\Phi(z)$ to the Fuchsian system (1) with coefficients being the matrices $B^{(i)}$, $i=1,\ldots,N$, is given by
(12) $\Phi=MD,$
where
(13) $D=\mathrm{diag}\left(\prod_{i=1}^{N}(z-a_{i})^{\beta_{1}^{(i)}},\ldots,\prod_{i=1}^{N}(z-a_{i})^{\beta_{p}^{(i)}}\right),$
and
(14) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{\kappa_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
where $q\vdash(l-k)$ represents the partitions of the integer $l-k$, and $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
Note that as $z$ varies, the contours $\gamma_{j}$ may be adjusted to avoid the points $P\in X_{a}$ with $\zeta(P)=z$ while staying in the same homology class from Theorem 4. In this way, starting from some $z_{0}\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$, a set of contours from Theorem 4, and the matrix $\Phi(z_{0})$ (12)-(14), we obtain $\Phi(z)$ for any $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ by analytical continuation of $\Phi(z_{0})$.
Example 5.
To illustrate the statement of the theorem we spell out the solutions in the matrix form for $p=5$. For $i=1,\dots,N,$ the $5\times 5$ matrices $B^{(i)}$ given by Theorem 1 for some constants $\theta_{i}\in\mathbb{R}$ is
(15) $B^{(i)}=\begin{pmatrix}\theta_{i}&\oint_{\gamma_{1}}\frac{w^{n}}{\zeta-a_{i}}\,\mathrm{d}\zeta&\oint_{\gamma_{2}}\frac{w^{2n}}{\zeta-a_{i}}\,\mathrm{d}\zeta&\oint_{\gamma_{3}}\frac{w^{3n}}{\zeta-a_{i}}\,\mathrm{d}\zeta&\oint_{\gamma_{4}}\frac{w^{4n}}{\zeta-a_{i}}\,\mathrm{d}\zeta\\ 0&\theta_{i}-\frac{n}{m}&\oint_{\gamma_{1}}\frac{w^{n}}{\zeta-a_{i}}\,\mathrm{d}\zeta&\oint_{\gamma_{2}}\frac{w^{2n}}{\zeta-a_{i}}\,\mathrm{d}\zeta&\oint_{\gamma_{3}}\frac{w^{3n}}{\zeta-a_{i}}\,\mathrm{d}\zeta\\ 0&0&\theta_{i}-\frac{2n}{m}&\oint_{\gamma_{1}}\frac{w^{n}}{\zeta-a_{i}}\,\mathrm{d}\zeta&\oint_{\gamma_{2}}\frac{w^{2n}}{\zeta-a_{i}}\,\mathrm{d}\zeta\\ 0&0&0&\theta_{i}-\frac{3n}{m}&\oint_{\gamma_{1}}\frac{w^{n}}{\zeta-a_{i}}\,\mathrm{d}\zeta\\ 0&0&0&0&\theta_{i}-\frac{4n}{m}\end{pmatrix}.$
Theorem 4 gives the following fundamental solution to the Fuchsian system (1) with the coefficients $B^{(i)}$ (15):
$\Phi(z)=MD,$
where
$M=\begin{pmatrix}1&\psi_{1}&\psi_{2}&\psi_{3}&\psi_{4}\\ 0&1&\psi_{1}&\psi_{2}&\psi_{3}\\ 0&0&1&\psi_{1}&\psi_{2}\\ 0&0&0&1&\psi_{1}\\ 0&0&0&0&1&\end{pmatrix}$
and
$D=\mathrm{diag}\left(\prod_{i=1}^{N}(z-a_{i})^{\theta_{i}},\prod_{i=1}^{N}(z-a_{i})^{\theta_{i}-\frac{n}{m}},\ldots,\prod_{i=1}^{N}(z-a_{i})^{\theta_{i}-\frac{4n}{m}}\right).$
The entries $\psi_{j}$ of the matrix $M$ are the following functions of $z$ expressed in terms of $\kappa_{j}$ given by Definition 2:
$\psi_{1}$ $=\kappa_{1},$
$\psi_{2}$ $=\kappa_{2}+\frac{1}{2}\kappa_{1}^{2},$
$\psi_{3}$ $=\kappa_{3}+\kappa_{1}\kappa_{2}+\frac{1}{6}\kappa_{1}^{3},$
$\psi_{4}$ $=\kappa_{4}+\kappa_{1}\kappa_{3}+\frac{1}{2}\kappa_{2}^{2}+\frac{1}{2}\kappa_{1}^{2}\kappa_{2}+\frac{1}{24}\kappa_{1}^{4}.$
The contour $\gamma_{j}$ for each $j=1,\dots,p-1$ in Theorem 4 is the same as the respective contour $\gamma_{j}$ in Theorem 1 used to construct the matrices $B^{(i)}$ serving as coefficients in (1). However, each set of contours $\gamma_{1},\dots,\gamma_{p-1}$ corresponding to a fixed solution $B^{(1)},\dots,B^{(N)}$ (7), (8) of the Schlesinger system and thus to a fixed Fuchsian system (1), gives rise to several different solutions of the Fuchsian system (corresponding to different initial conditions for (1)). This happens because, as stated in Theorem 4, the contours $\gamma_{j}$ are elements of the homology of the surface punctured at the points $P\in X_{a}$ for which $\zeta(P)=z.$ In other words, a given contour $\gamma_{j}$ in $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s},P_{a_{1}},\ldots,P_{a_{N}}\},\mathbb{C})$ will give different values of the integral $\kappa_{j}$ (9) depending on the positions of the points of the surface satisfying $\zeta(P)=z_{0}$ relative to the contour $\gamma_{j}.$ Here we assume that initial value for $\Phi$ is given at $z=z_{0}$. Let us illustrate this with a simple example.
Example 6.
Let all contours $\gamma_{j}$ for $j=1,\dots,p-1$ be equal to the same trivial element $\gamma$ of $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s},P_{a_{1}},\ldots,P_{a_{N}}\},\mathbb{C})$. In this case, the matrices $B^{(i)}$ given by Theorem 1 are diagonal:
(16) $B^{(i)}={\rm diag}\left(\theta_{i},\theta_{i}-\frac{n}{m},\theta_{i}-\frac{2n}{m},\dots,\theta_{i}-\frac{(p-1)n}{m}\right).$
Consider now the points $P\in X_{a}$ for which $\zeta(P)=z$ for some $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$. There are $m$ such points and we denote them by $P_{z}^{(1)},\dots,P_{z}^{(m)}.$ The integrals $\kappa_{j}$ (9) computed with $\gamma_{j}=\gamma$ depend on how many of the points $P_{z}^{(k)}$ belong to the interior of $\gamma$, that is to the domain delimited by the contour $\gamma$ homologous to zero on $X_{a}$.
Let us introduce the short hand notation
(17) $\Pi=\prod_{i=1}^{N}(z-a_{i})\qquad\mbox{and}\qquad\Pi^{\theta}=\prod_{i=1}^{N}(z-a_{i})^{\theta_{i}}$
for $\theta:=(\theta_{1},\dots,\theta_{p})\in\mathbb{C}^{p}$ and consider two cases:
  • Case 1:
    The contour $\gamma$ does not enclose any of the points $P_{z}^{(k)}$. In other words, $\gamma$ is a trivial element of $H_{1}(X_{a}\setminus(\{P_{a_{1}},\ldots,P_{a_{N}},\infty_{1},\dots,\infty_{s}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z\}),{\mathbb{C}})$. In this case each of the $\kappa_{j}$ with $\gamma_{j}=\gamma$ vanishes and the solution $\Phi(z)$ given by Theorem 4 is diagonal: $M$ (14) is the identity matrix and $\Phi(z)=D:$
    (18) $\Phi(z)=\mathrm{diag}\left(\Pi^{\theta},\Pi^{\theta-\frac{n}{m}},\ldots,\Pi^{\theta-\frac{(p-1)n}{m}}\right),$
    where we use $\theta-\frac{n}{m}:=(\theta_{1}-\frac{n}{m},\dots,\theta_{p}-\frac{n}{m}).$
  • Case 2:
    The contour $\gamma$ contains $t<m$ points $P_{z}^{(1)},\dots,P_{z}^{(t)}$ in its interior. In this case, we have for $\kappa_{j}$
    (19) $\kappa_{j}:=-\frac{m}{jn}\oint_{\gamma}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\underset{P_{z}^{(r)}}{\rm res}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\epsilon_{r}\Pi^{\frac{jn}{m}},$
    where $\epsilon_{r}$ are different $m$th roots of unity. Let us, for simplicity of notation, denote $\alpha_{j}:=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\epsilon_{r}$ so that $\kappa_{j}=\alpha_{j}\Pi^{\frac{jn}{m}}.$ Now, for the entries above the diagonal of the matrix $M$ given by (14), we have:
    $M_{kl}=\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{\left(\alpha_{j}\Pi^{\frac{jn}{m}}\right)^{\sigma_{q}(j)}}{\sigma_{q}(j)!}=\hat{\alpha}_{l-k}\Pi^{\frac{(l-k)n}{m}}$
[…]
    (20) $\Phi(z)=\begin{pmatrix}\Pi^{\theta}&\hat{\alpha}_{1}\Pi^{\theta}&\hat{\alpha}_{2}\Pi^{\theta}&\dots&\hat{\alpha}_{p-1}\Pi^{\theta}\\ 0&\Pi^{\theta-\frac{n}{m}}&\hat{\alpha}_{1}\Pi^{\theta-\frac{n}{m}}&\dots&\hat{\alpha}_{p-2}\Pi^{\theta-\frac{n}{m}}\\ 0&0&\Pi^{\theta-\frac{2n}{m}}&\dots&\hat{\alpha}_{p-3}\Pi^{\theta-\frac{2n}{m}}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ 0&0&0&0&\Pi^{\theta-\frac{(p-1)n}{m}}&\end{pmatrix},$
    or $\Phi_{kl}=\hat{\alpha}_{l-k}\Pi^{\theta-\frac{k-1}{m}}$ for $l\geqslant k$ and $\Phi_{kl}=0$ otherwise.
We thus obtain two solutions (18) and (20) for the same system (1) with coefficients (16).
To prepare for the proof of Theorem 4, we first establish the following lemma.
Lemma 7.
Let $M$ and $D$ be the upper-triangular $p\times p$ matrices defined in Theorem 4 and $\kappa^{\prime}_{r}$ be the derivative (10) of $\kappa_{r}$ with respect to $z$. Then,
  1. (1)
    $\frac{\mathrm{d}}{\mathrm{d}z}D=D_{z}=\mathrm{diag}\left(\sum_{i=1}^{N}\frac{\beta_{1}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{1}^{(i)}},\ldots,\sum_{i=1}^{N}\frac{\beta_{p}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{p}^{(i)}}\right)$,
  2. (2)
    $\frac{\mathrm{d}}{\mathrm{d}z}M=M_{z}$ for which the $(kl)$-entry is
    $(M_{z})_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\sum_{r=1}^{l-k}\left(\frac{\kappa_{r}^{\sigma_{q}(r)-1}}{(\sigma_{q}(r)-1)!}\kappa_{r}^{\prime}\prod_{j=1,j\neq r}^{l-k}\frac{\kappa_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\right)\quad&\text{when } $l>k$,\\ \quad 0\quad&\text{when } $l\leq k$,\end{dcases*}$
  3. (3)
    $D^{-1}=\mathrm{diag}\left(\prod_{j=1}^{N}(z-a_{j})^{-\beta_{1}^{(i)}},\ldots,\prod_{j=1}^{N}(z-a_{j})^{-\beta_{p}^{(i)}}\right)$, and
  4. (4)
    the inverse of the matrix $M$ has the following entries
    (21) $(M^{-1})_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{(-\kappa_{j})^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$.\end{dcases*}$
Proof.   Items (1) - (3) of the lemma being obtained by straightforward observation, let us prove item (4). We want to show that
(22) $(MM^{-1})_{kl}=\delta_{lk}$
with $M^{-1}$ given in the statement of the proposition. Since both $M$ (14) and $M^{-1}$ (21) are upper triangular with 1’s on the main diagonal, the same holds for their product: $(MM^{-1})_{kl}=\delta_{lk}$ for $l\leqslant k.$ Let us consider the case when $l>k$. Using that $M_{kn}=(M^{-1})_{kn}=0$ for $k>n$, we have
$(MM^{-1})_{kl}$ $=\sum_{n=1}^{p}M_{kn}(M^{-1})_{nl}=\sum_{n=k}^{l}M_{kn}(M^{-1})_{nl}$
$=\sum_{n=k}^{l}\left[\left(\sum_{r_{1}\,\vdash\,(n-k)}\,\prod_{j_{1}=1}^{n-k}\frac{\kappa_{j_{1}}^{\sigma_{r_{1}}(j_{1})}}{\sigma_{r_{1}}(j_{1})!}\right)\left(\sum_{r_{2}\,\vdash\,(l-n)}\,\prod_{j_{2}=1}^{l-n}\frac{(-\kappa_{j_{2}})^{\sigma_{r_{2}}(j_{2})}}{\sigma_{r_{2}}(j_{2})!}\right)\right].$
Let us show that this expression is zero. Let $T$ be a strictly positive integer and $K,L\geqslant 0.$ Using that $\sigma_{q}(j)=0$ for $q\vdash T$ and $j>T$, note that
$\sum_{K+L=T}\left(\sum_{r_{1}\vdash K}\prod_{j=1}^{K}\frac{\kappa_{j}^{\sigma_{r_{1}}(j)}}{\sigma_{r_{1}}(j)!}\right)\left(\sum_{r_{2}\vdash L}\prod_{j=1}^{L}\frac{(-\kappa_{j})^{\sigma_{r_{2}}(j)}}{\sigma_{r_{2}}(j)!}\right)=\sum_{K+L=T}\sum_{r_{1}\vdash K}\sum_{r_{2}\vdash L}\prod_{j=1}^{T}\frac{\kappa_{j}^{\sigma_{r_{1}}(j)}(-\kappa_{j})^{\sigma_{r_{2}}(j)}}{\sigma_{r_{1}}(j)!\sigma_{r_{2}}(j)!}\\ =\sum_{q\vdash T}\sum_{\begin{subarray}{c}(r_{1},r_{2})\\ r_{1}+r_{2}=q\end{subarray}}\prod_{j=1}^{T}\frac{\kappa_{j}^{\sigma_{r_{1}}(j)}(-\kappa_{j})^{\sigma_{r_{2}}(j)}}{\sigma_{r_{1}}(j)!\sigma_{r_{2}}(j)!}.$
Here $r_{1}+r_{2}$ stands for the sum of two partitions combining the parts of $r_{1}$ and $r_{2}$, such that $\sigma_{r_{1}}(j)+\sigma_{r_{2}}(j)=\sigma_{r_{1}+r_{2}}(j)$ for any $j$ and for two partitions $q_{1}$ and $q_{2}$ the pairs $(q_{1},q_{2})$ and $(q_{2},q_{1})$ are considered different. Since the numbers $\sigma_{q}(1),\dots\sigma_{q}(T)$ define the partition $q$ completely, we have for every fixed partition $q$
(23) $\prod_{j=1}^{T}\sum_{\sigma_{r_{1}}+\sigma_{r_{2}}=\sigma_{q}(j)}\frac{\kappa_{j}^{\sigma_{r_{1}}}(-\kappa_{j})^{\sigma_{r_{2}}}}{\sigma_{r_{1}}!\sigma_{r_{2}}!}=\sum_{\begin{subarray}{c}(r_{1},r_{2})\\ r_{1}+r_{2}=q\end{subarray}}\prod_{j=1}^{T}\frac{\kappa_{j}^{\sigma_{r_{1}}(j)}(-\kappa_{j})^{\sigma_{r_{2}}(j)}}{\sigma_{r_{1}}(j)!\sigma_{r_{2}}(j)!}.$
Thus, continuing our calculation, we obtain
$\sum_{K+L=T}\left(\sum_{r_{1}\vdash K}\prod_{j=1}^{K}\frac{\kappa_{j}^{\sigma_{r_{1}}(j)}}{\sigma_{r_{1}}(j)!}\right)\left(\sum_{r_{2}\vdash L}\prod_{j=1}^{L}\frac{(-\kappa_{j})^{\sigma_{r_{2}}(j)}}{\sigma_{r_{2}}(j)!}\right)\\ =\sum_{q\vdash T}\prod_{j=1}^{T}\frac{1}{\sigma_{q}(j)!}\sum_{\sigma_{r_{1}}+\sigma_{r_{2}}=\sigma_{q}(j)}\frac{\sigma_{q}(j)!}{\sigma_{r_{1}}!\sigma_{r_{2}}!}\kappa_{j}^{\sigma_{r_{1}}}(-\kappa_{j})^{\sigma_{r_{2}}}=\sum_{q\vdash T}\prod_{j=1}^{T}\frac{1}{\sigma_{q}(j)!}(\kappa_{j}-\kappa_{j})^{\sigma_{q}(j)}=0.$
For $T=l-k$, this implies (22) with $l>k$ and thus proves that (21) defines the inverse of the matrix $M$ (14). $\Box$ We now have all the necessary tools to prove Theorem 4.
Proof of Theorem 4. We want to show that $\Phi=MD$ where both $M$ and $D$ are defined in Theorem 4, is a fundamental solution of the Fuchsian system (1), i.e.
$\Phi_{z}\Phi^{-1}=\sum_{i=1}^{N}\frac{1}{z-a_{i}}B^{(i)},$
where $\Phi_{z}=\frac{\mathrm{d}}{\mathrm{d}z}\Phi.$ Differentiating (12), we have
(24) $\Phi_{z}\Phi^{-1}=(M_{z}D+MD_{z})D^{-1}M^{-1}=M_{z}M^{-1}+MD_{z}D^{-1}M^{-1}.$
The matrix $M_{z}M^{-1}$ is upper triangular as a product of upper triangular matrices. Its diagonal entries are zeros since this is the case for $M_{z}$, see Lemma 7. Let us consider the entries of $M_{z}M^{-1}$ lying above the main diagonal. Fix $k<l$ and set $j=l-k$. By Lemma 7, for $r\geqslant 1$, we have
$(M_{z})_{k,k+r}=\sum_{q_{1}\vdash r}\sum_{t=1}^{r}\frac{\kappa_{t}^{\prime}}{\kappa_{t}}\;\sigma_{q_{1}}(t)\prod_{i=1}^{r}\frac{\kappa_{i}^{\sigma_{q_{1}}(i)}}{\sigma_{q_{1}}(i)!}$
and for $1\leqslant r<j$
(25) $(M^{-1})_{k+r,k+j}=\sum_{q_{2}\vdash(j-r)}\prod_{i=1}^{j-r}\frac{(-\kappa_{i})^{\sigma_{q_{2}}(i)}}{\sigma_{q_{2}}(i)!},$
hence
$(M_{z}M^{-1})_{kl}=\sum_{r=0}^{j}\;\sum_{q_{1}\vdash r}\;\sum_{q_{2}\vdash(j-r)}\;\sum_{t=1}^{r}\frac{\kappa_{t}^{\prime}}{\kappa_{t}}\,\sigma_{q_{1}}(t)\left(\prod_{i=1}^{r}\frac{\kappa_{i}^{\sigma_{q_{1}}(i)}}{\sigma_{q_{1}}(i)!}\right)\;\prod_{h=1}^{j-r}\frac{(-\kappa_{h})^{\sigma_{q_{2}}(h)}}{\sigma_{q_{2}}(h)!}.$
In this proof, we assume that there exists one partition of zero. Grouping the partitions $q_{1}$ and $q_{2}$ into a sum, we obtain the sum over all partitions $q:=q_{1}+q_{2}$ of $j$ with $\sigma_{q_{1}}(i)+\sigma_{q_{2}}(i)=\sigma_{q}(i)$. This gives for $j=l-k\geqslant 1$
$(M_{z}M^{-1})_{kl}=\sum_{q\vdash j}\;\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\vdash j\end{subarray}}\;\sum_{t=1}^{j}\frac{\kappa_{t}^{\prime}}{\kappa_{t}}\,\sigma_{q_{1}}(t)\;\prod_{i=1}^{j}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}(-1)^{\sigma_{q_{2}}(i)}\frac{\kappa_{i}^{\sigma_{q}(i)}}{\sigma_{q}(i)!}$
or, equivalently,
(26) $(M_{z}M^{-1})_{kl}=\sum_{q\vdash j}\;\left(\prod_{i=1}^{j}\frac{\left(\kappa_{i}\right)^{\sigma_{q}(i)}}{\sigma_{q}(i)!}\right)\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\vdash j\end{subarray}}\sum_{t=1}^{j}\frac{\kappa_{t}^{\prime}}{\kappa_{t}}\,\sigma_{q_{1}}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{q_{2}}(i)}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}.$
Let us now consider the contribution, denoted by $C(q,t),$ to (26) of a fixed partition $q\vdash j$ and of a fixed value of $t$. Introduce the vector $\sigma_{q}=(\sigma_{q}(1),\dots,\sigma(j))$, which defines the partition $q$ uniquely. We can rewrite the contribution of $q$ and $t$ in (26) as a sum over all decompositions of $\sigma_{q}$ into the sum of two vectors. Then, similarly to (23), this becomes the product of sums (we write $\sigma_{k}(i)$ for the $i$th component of the vector $\sigma_{k}$):
$C(q,t)=\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\end{subarray}}\sigma_{q_{1}}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{q_{2}}(i)}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}=\sum_{\sigma_{1}+\sigma_{2}=\sigma_{q}}\sigma_{1}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{2}(i)}\binom{\sigma_{q}(i)}{\sigma_{1}(i)}$
(27) $=\left(\sum_{x_{t}=0}^{\sigma_{q}(t)}x_{t}\binom{\sigma_{q}(t)}{x_{t}}(-1)^{\sigma_{q}(t)-x_{t}}\right)\prod_{\begin{subarray}{c}i=1\\ i\neq t\end{subarray}}^{j}\left(\sum_{x_{i}=0}^{\sigma_{q}(i)}\binom{\sigma_{q}(i)}{x_{i}}(-1)^{\sigma_{q}(i)-x_{i}}\right).$
If $\sigma_{q}(i)\neq 0$, the sum over $x_{i}$ for $i\neq t$ equals $(1+x)^{\sigma_{q}(i)}$ evaluated at $x=-1$ and thus vanishes. If $\sigma_{q}(i)=0$, the sum over $x_{i}$ is equal to $1$. The sum over $x_{t}$, if $\sigma_{q}(t)>1$, can be rewritten as the derivative of $(1+x)^{\sigma_{q}(t)}$ evaluated at $-1$, therefore it is equal to $1$ if $\sigma_{q}(t)=1$ and 00 otherwise. Thus, $C(q,t)$ is non-zero if and only if $q=j$ and $t=j$, that is if $q$ is the partition of $j$ containing one part only, in which case $C(q,t)=1$. Using this result in (26), and noting that the only splitting of $j$ into the sum of two partitions that gives a non-zero contribution in (26) is $q_{1}=j,\;q_{2}=0$, we obtain
$(M_{z}M^{-1})_{kl}=\begin{cases}0,&l\leq k,\\ \kappa_{\,l-k}^{\prime},&l>k.\end{cases}$
Let us now consider the term $MD_{z}D^{-1}M^{-1}$ in (24). Recalling that $\beta_{h}^{(i)}=\beta_{k}^{(i)}-\tfrac{(h-k)n}{m}$ and using Lemma 7 for $D_{z}$, one finds that for $l\geqslant k$
$(MD_{z}D^{-1}M^{-1})_{kl}$ $=\sum_{i=1}^{N}\sum_{h=k}^{l}M_{kh}\,\frac{\beta_{h}^{(i)}}{\;z-a_{i}\!}\,(M^{-1})_{hl}$
$=\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}}\!\sum_{h=k}^{l}M_{kh}(M^{-1})_{hl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{h=k}^{l}(h-k)M_{kh}(M^{-1})_{hl}$
$=\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}}\delta_{kl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{r=0}^{l-k}rM_{k,k+r}(M^{-1})_{k+r,l}.$
Comparing with our above derivation, we see that the second term can be rewritten using partitions similarly to (26), where $j=l-k$. The factor of $r$ corresponds to the sum of all the parts of the partition $q_{1}$. More precisely, using (14) for $M$ as well as formulas from (25) to (3), we obtain
(28) $\sum_{r=0}^{l-k}rM_{k,k+r}(M^{-1})_{k+r,l}$ $=\sum_{q\vdash j}\;\left(\prod_{i=1}^{j}\frac{\left(\kappa_{i}\right)^{\sigma_{q}(i)}}{\sigma_{q}(i)!}\right)\!\!\!\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\vdash j\end{subarray}}\!\!\!\sum_{t=1}^{j}t\,\sigma_{q_{1}}(t)\prod_{i=1}^{j}(-1)^{\sigma_{q_{2}}(i)}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}$
$=\sum_{q\vdash j}\left(\prod_{i=1}^{j}\frac{\left(\kappa_{i}\right)^{\sigma_{q}(i)}}{\sigma_{q}(i)!}\right)\sum_{t=1}^{j}t\,C(q,t).$
Given that $C(q,t)$ is non-zero only in the case of one term partition $q=j=l-k$ and $t=j,$ in which case $C(j,j)=1$, we obtain that quantity (28) is equal to $j\kappa_{j}$ and thus
$(MD_{z}D^{-1}M^{-1})_{kl}=\begin{cases}0,&l<k,\\ \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}},&l=k,\\ \displaystyle-\,\frac{(l-k)n}{m}\,\kappa_{\,l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
Summing the two contributions into (24), we have
$(\Phi_{z}\Phi^{-1})_{kl}=\begin{cases}0,&l<k,\\[3.0pt] \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}},&l=k,\\[4.0pt] \kappa_{l-k}^{\prime}-\displaystyle\frac{({l-k})n}{m}\,\kappa_{l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
Recalling that from Lemma 3, for each $l-k\geq 1$,
$\;\kappa_{l-k}^{\prime}-\frac{({l-k})n}{m}\kappa_{l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}}=\sum_{i=1}^{N}\frac{B^{(i)}_{k,l}}{z-a_{i}},$
we obtain
$\Phi_{z}\Phi^{-1}=\sum_{i=1}^{N}\frac{B^{(i)}}{z-a_{i}}.$
$\Box$
4.  Monodromy matrices of solutions
Let us introduce a basis $\{\mathcal{A}_{1},\dots,\mathcal{A}_{g};\mathcal{B}_{1},\dots,\mathcal{B}_{g}\}$ in $H_{1}(X_{a},\mathbb{C})$ not passing trough ramification points of the covering $\zeta$ with the intersection indices being $\mathcal{A}_{i}\circ\mathcal{A}_{k}=\mathcal{B}_{i}\circ\mathcal{B}_{k}=0$ and $\mathcal{A}_{i}\circ\mathcal{B}_{k}=1$; such a basis is called canonical. The integration contours $\gamma_{j}$ from Theorem 4 are linear combinations of the cycles $\mathcal{A}_{k},\mathcal{B}_{k}$ for $k=1,\dots,g$, small loops $\gamma_{a_{i}}$ encircling the points $P_{a_{i}}$ for $i=1,\dots,N$, small loops $\gamma_{\infty_{\alpha}}$ encircling the points $\infty_{\alpha}$ for $\alpha=1,\dots,s$ as well as small loops $\eta_{t}$ for $t=1,\dots,m$ each encircling a point $P_{z}^{(t)}$ for which $\zeta(P)=z.$ Assume that the only pairs of contours that intersect each other are $\mathcal{A}_{k}$ and $\mathcal{B}_{k}$ for $k=1,\dots,g.$
Let us bring all the contours $\{\mathcal{A}_{k},\mathcal{B}_{k}\}_{k=1}^{g}$ to pass through one point $R_{0}\in X_{a}$ such that $\zeta(R_{0})=z_{0}$ for some choice of $z_{0}\in\mathbb{C}\setminus\{a_{1},\ldots,a_{N}\}$ and let us identify the $z$- and $\zeta$-spheres. Now, we can choose generators $\rho_{1},\dots,\rho_{N}$ of the fundamental group $\pi_{1}(\mathbb{C}\setminus\{a_{1},\dots,a_{N}\},z_{0})$ in a way so that they do not intersect the projections $\zeta(\mathcal{A}_{k}),\zeta(\mathcal{B}_{k})$, $\zeta(\gamma_{a_{i}})$ and $\zeta(\gamma_{\infty_{\alpha}})$ of the above contours (besides at the point $z_{0}$ where the intersection is not transversal). We assume also that $\rho_{j}$ goes anticlockwise once around the point $z=a_{j}$ and does not encircle any other branch point of the covering $\zeta.$
For this configuration of the contours we may analytically continue the solution $\Phi(z_{0})$ to the Fuchsian system (1) along the generators $\rho_{i}$ to obtain another solution $\Phi_{\rho_{i}}(z_{0})$ of the same system. The system being linear, the two solutions are related by $\Phi_{\rho_{i}}(z_{0})=\Phi(z_{0})\mathcal{M}_{i}$ for some (monodromy) matrix $\mathcal{M}_{i}$. Given that the coefficients $B^{(i)}$ of (1) solve the Schlesinger system, the monodromy matrices are independent of small variations of $a_{1},\dots,a_{N}.$
In this section, we calculate the monodromy matrices for three sets of contours $\gamma_{1},\dots,\gamma_{p-1}$. The general case can be reconstructed from the three cases we consider here.
Case 1.
Assume that each of the contours $\gamma_{1},\dots,\gamma_{p-1}$ is a linear combination of the contours $\mathcal{A}_{k},\mathcal{B}_{k}$, $\gamma_{a_{i}}$ and $\gamma_{\infty_{\alpha}}$ for $k=1,\dots,g$, $i=1,\dots,N$, and $\alpha=1,\dots,s$. In other words, we assume that the contours $\eta_{t}$ for $t=1,\dots,m$ are not present in the linear combinations. In this case, the functions $\kappa_{j}(z)$ given by (9) do not change under the analytical continuation along the loops $\rho_{j}$ since the contours $\gamma_{j}$ are not affected by this operation. Therefore, the matrix $M$ (14) does not change and the only monodromy of the solution $\Phi(z_{0})$ comes from the monodromy of the diagonal matrix $D$ (13). Thus we obtain that
$\mathcal{M}_{i}=\mathrm{diag}\left(e^{2\pi\mathrm{i}\beta_{1}^{(i)}},\ldots,e^{2\pi\mathrm{i}\beta_{p}^{(i)}}\right)=\mathrm{diag}\left(e^{2\pi\mathrm{i}\theta_{i}},e^{2\pi\mathrm{i}\theta_{i}-\frac{n}{m}},\ldots,e^{2\pi\mathrm{i}\theta_{i}-\frac{(p-1)n}{m}}\right).$
Case 2.
Assume now that all the contours $\gamma_{j}$ are obtained as linear combinations of the contours $\eta_{1},\dots,\eta_{m}$ each encircling a point of the set $\{P_{z}^{(1)},\dots,P_{z}^{(m)}\}$. In this case, the functions $\kappa_{j}(z)$ are linear combinations of residues at the points $P_{z}^{(t)}$ and thus are calculated in (19). The solution $\Phi(z_{0})$ has the form (20) with $z=z_{0}$, some constants $\hat{\alpha}_{j}$ (possibly different from those in (20) to correct for different contours $\gamma_{j}$), and $\Pi^{\theta}$ given by (17). We give the monodromy matrices for the case $p=3$:
$\mathcal{M}_{i}={\rm e}^{2\pi\bf{\rm i}\theta_{i}}\begin{pmatrix}1&\hat{\alpha}_{1}\left(1-{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)&\hat{\alpha}_{2}\left(1-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)-\hat{\alpha}_{1}^{2}\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}&\hat{\alpha}_{1}\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\end{pmatrix}.$
Case 3.
Finally, we consider an example where some of the contours $\gamma_{j}$ go around points $P_{z}^{(t)}$ and some other are like in Case 1 above. Let us again consider the case $p=3$ and set $\gamma_{1}=\sum_{r=1}^{m}c_{r}\eta_{r}$ for some $c_{r}\in\mathbb{C}$ and let $\gamma_{2}$ to be one of the contours from Case 1. Then $\kappa_{2}(z)$ does not transform under the analytical continuation along the loops $\rho_{i}$ while $\kappa_{1}=\alpha\Pi^{\frac{n}{m}}$ as shown in Example 6 for some $\alpha\in\mathbb{C}$ and thus does transform. The solution given by Theorem 4 is
$\Phi(z)=\begin{pmatrix}\Pi^{\theta}&\kappa_{1}\Pi^{\theta-\frac{n}{m}}&\left(\kappa_{2}+\frac{1}{2}\kappa_{1}^{2}\right)\Pi^{\theta-\frac{2n}{m}}\\ 0&\Pi^{\theta-\frac{n}{m}}&\kappa_{1}\Pi^{\theta-\frac{2n}{m}}\\ 0&0&\Pi^{\theta-\frac{2n}{m}}\end{pmatrix}=\Pi^{\theta}\begin{pmatrix}1&\alpha&\kappa_{2}\Pi^{-\frac{2n}{m}}+\frac{1}{2}\alpha^{2}\\ 0&\Pi^{-\frac{n}{m}}&\alpha\Pi^{-\frac{n}{m}}\\ 0&0&\Pi^{-\frac{2n}{m}}\end{pmatrix}.$
The monodromy matrices are for $i=1,\dots,N$
$\mathcal{M}_{i}={\rm e}^{2\pi\bf{\rm i}\theta_{i}}\begin{pmatrix}1&\alpha\left(1-{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)&\frac{1}{2}\alpha^{2}\left(1+{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}-2{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}&\alpha\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\end{pmatrix}.$
We see that in all cases the monodromy matrices are independent of small variations of $a_{i}$, $i=1,\dots,N$, and are conjugate to ${\rm exp}\{2\pi\bf{\rm i}\,B^{(i)}\}$ as expected according to the general theory of Fuchsian systems, see for example, [16]. Note that our solutions also transform when $z$ goes around the point at infinity, and the corresponding monodromy matrix $\mathcal{M}_{\infty}$ is conjugate to ${\rm exp}\{2\pi\bf{\rm i}\,B^{(\infty)}\}$ where $B^{(\infty)}=\sum_{i=1}^{N}B^{(i)}$ is a nonzero diagonal matrix, see the end of Section 2. A loop around $z=\infty$ is given by a product of loops $\rho_{i}$ and thus $\mathcal{M}_{\infty}$ is obtained as the corresponding product of the matrices $\mathcal{M}_{i}$.
5.  Some rational solutions
As mentioned after Theorem 1, the contours of integration may be chosen in a way for the solutions of the Schlesinger system from Theorem 1 to be given by residues of the differentials (6). From [7], we know that for $n>0$, the differentials only have poles at the $s$ points at infinity $\infty_{1},\dots,\infty_{s}$. For coprime $m$ and $N$, there is only one pole at infinity and thus the residue there is zero. For $n<0$, all poles are at the finite ramification points $P_{a_{1}},\dots,P_{a_{N}}.$ In these cases we obtain rational solutions to the Schlesinger system and to the Fuchsian system as well.
5.1  Positive values of $n$
Following [7], let us denote $N=sN_{1}\,,$ $m=sm_{1}\,$ for $s={\rm gcd}(m,N)$ and recall that $n>0$ should be coprime with $m$. In this case the differential $\Omega_{i}^{(j)}(a)$ (6) has $s$ poles, one at each of the points $\infty_{1},\ldots,\infty_{s}$ at infinity of the Riemann surface $X_{a}$. The behaviour of these differentials at $P=(\zeta,w)\sim\infty_{\alpha}$ for $\alpha=1,\dots,s$ in terms of the local parameter $\xi:=\xi_{\infty_{\alpha}}$ given by (2) is
(29) $\Omega_{i}^{(j)}(P)=\frac{w^{jn}d\zeta}{\zeta-a_{i}}=\frac{\nu_{jn\alpha}\prod_{i=1}^{N}(1-a_{i}\xi^{m_{1}})^{jn/m}}{\xi^{jnN_{1}+1}(1-a_{i}\xi^{m_{1}})}\,d\xi,\qquad\nu_{jn\alpha}=-m_{1}\,e^{2\pi\bf{\rm i}jn(\alpha-1)/s}.$
We thus have a pole of order $jnN_{1}+1$ at $P=\infty_{\alpha}$ for each $\Omega_{i}^{(j)}$. As can be seen from (29) the residue at the pole is nonzero only if $j$ is a multiple of $m_{1}$.
Note that the constant overall factor in (29) was unimportant in [7] but for our purposes in the present paper, we need to keep track of it. To this end, it is convenient introduce the following notation
(30) $A_{j}:=(-1)^{N\frac{jn}{m}}\nu_{jn\alpha}.$
Let $\gamma_{\infty_{\alpha}}$ be a small contour encircling $\infty_{\alpha}$. Choosing now $\gamma_{l-k}=\frac{c_{l-k}}{2\pi\bf{\rm i}A_{l-k}}\gamma_{\infty_{\alpha}}$ for some $c_{l-k}\in\mathbb{C}$, we have a polynomial solution to the Schlesinger system given, for some $\theta_{i}\in\mathbb{C}$, by (see Theorem 3 of [7])
(31) $B^{(i)}_{kl}=\begin{dcases*}\frac{c_{l-k}}{A_{l-k}}\;\underset{\infty_{\alpha}}{\rm res}\,\Omega_{i}^{(l-k)}(a),\quad\text{if }(l-k)/m_{1}\in\mathbb{Z}\;\text{ and }\;l>k\\ 0,\qquad\qquad\qquad\qquad\;\;\,\text{if }(l-k)/m_{1}\notin\mathbb{Z}\;\text{ or }\;l<k\\ \theta_{i}-\frac{(k-1)n}{m},\qquad\quad\;\;\,\text{if }\;l=k.\end{dcases*}$
The nonzero residue from (31), for $(l-k)/m_{1}\in\mathbb{Z}$, is explicitly given by [7]
(32) $\underset{\infty_{\alpha}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=A_{l-k}\sum_{k_{1}+\ldots+k_{N}+q=N(l-k)\frac{n}{m}}(-1)^{q}{(l-k)\frac{n}{m}\choose k_{1}}\ldots{(l-k)\frac{n}{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}a_{i}^{q},$
with the generalized binomial coefficients
(${\beta\choose j}=\frac{\beta(\beta-1)\cdots(\beta-j+1)}{j!},\qquad{\beta\choose 0}=1$
defined for $\beta\in\mathbb{R}$ and $j\in\mathbb{N}$. Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (31) are rational functions of $z$ given by the next corollary.
Corollary 8.
Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and $p,n,m,N$ be strictly positive integers such that ${\rm gcd}(n,m)=1$ and $s={\rm gcd}(m,N)>1$. With $c_{j}\in\mathbb{C}$ as in (31), let $R_{j}=R_{j}(z)$ be the rational funtion
(33) $R_{j}(z)=\begin{dcases*}-\frac{m\,c_{j}}{jn}\sum_{k_{1}+\ldots+k_{N}+q=N\frac{jn}{m}}(-1)^{q}{{jn}/{m}\choose k_{1}}\ldots{{jn}/{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{q},\quad\text{if }j/m_{1}\in\mathbb{Z}\\ 0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\;\text{if }j/m_{1}\notin\mathbb{Z}.\end{dcases*}$
Denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$ and $i=1,\dots,N$. Then the following upper triangular matrix $\Phi(z)$ solves the Fuchsian system (1) with coefficients $B^{(i)}$ being the upper triangular $p\times p$ matrices given by (31).
(34) $\Phi=MD,$
where
(35) $D=\mathrm{diag}\left(\prod_{i=1}^{N}(z-a_{i})^{\beta_{1}^{(i)}},\ldots,\prod_{i=1}^{N}(z-a_{i})^{\beta_{p}^{(i)}}\right),$
and
(36) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{R_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
where $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
The $p=2$ and $\theta_{i}=\frac{n}{2m}$ case of this theorem obtained in Theorem 1 of [1] was the starting point of the present paper.
Proof.   Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$. For $n>0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $\infty_{1},\dots,\infty_{s}$ as explained at the beginning of Section 5.1. Comparing the statements of Corollary 8 and Theorem 4, we only need to prove that $\kappa_{j}$ from Definition 2 coincides with $R_{j}$ (33) for the choice of the integration contour $\gamma_{j}=\frac{c_{j}}{2\pi\bf{\rm i}A_{j}}\gamma_{\infty_{\alpha}}$ with $\gamma_{\infty_{\alpha}}$ being a small contour encircling $\infty_{\alpha}\in X_{a}$ and not containing in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$. With this choice and $m_{1},N_{1}$ as above, we have
$\kappa_{j}=-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\infty_{\alpha}}{\rm res}\,\frac{w^{jn}\mathrm{d}\zeta}{\zeta-z}=-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}\prod_{i=1}^{N}(1-a_{i}\xi^{m_{1}})^{jn/m}}{\xi^{jnN_{1}+1}(1-z\xi^{m_{1}})}\,d\xi.$
where $\xi:=\xi_{\infty_{\alpha}}$ is the local parameter at $\infty_{\alpha}$ given by (2) and $\nu_{jn\alpha}$ as in (29). Therefore, expanding every factor in Taylor series in $\xi\sim 0$,
$\kappa_{j}$ $=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{k_{1}=0}^{\infty}{jn/m\choose k_{1}}(-a_{1}\xi^{m_{1}})^{k_{1}}\ldots\sum_{k_{N}=0}^{\infty}{jn/m\choose k_{N}}(-a_{N}t^{m_{1}})^{k_{N}}\sum_{q=0}^{\infty}(z\xi^{m_{1}})^{q}$
$=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{r=0}^{\infty}\Bigl{[}\sum_{k_{1}+\ldots+k_{N}+q=r}(-1)^{r-q}{jn/m\choose k_{1}}\ldots{jn/m\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{q}\Bigr{]}\,\xi^{rm_{1}}.$
To obtain the residue we need to take the term of the series corresponding to the value of $r$ such that $jnN_{1}=rm_{1}$. Note that the constant $A_{j}$ cancels out and we obtain $R_{j}$ (33). $\Box$
Example 9.
Let us set, for simplicity, $\theta_{1}=\theta_{2}=\frac{1}{4}$, $p=2$, and $n=1,\;m=2,\;N=2$ so that $s=2$. In this case, $\beta_{1}=\frac{1}{4}$ and $\beta_{2}=-\frac{1}{4}$ and the polynomial in $a_{1},\,a_{2}$ solution (31) to the Schlesinger system is
$B^{(1)}=\begin{pmatrix}\frac{1}{4}&\frac{c_{1}}{2}(a_{2}-a_{1})\\ 0&-\frac{1}{4}\end{pmatrix},\qquad B^{(2)}=\begin{pmatrix}\frac{1}{4}&\frac{c_{1}}{2}(a_{1}-a_{2})\\ 0&-\frac{1}{4}\end{pmatrix}.$
The solution given by Corollary 8 to the corresponding Fuchsian system is
$\Phi(z)=\begin{pmatrix}(z-a_{1})^{\frac{1}{4}}(z-a_{2})^{\frac{1}{4}}&\frac{c_{1}(2z-a_{1}-a_{2})}{(z-a_{1})^{\frac{1}{4}}(z-a_{2})^{\frac{1}{4}}}\\ \\ 0&\frac{1}{(z-a_{1})^{\frac{1}{4}}(z-a_{2})^{\frac{1}{4}}}\end{pmatrix}.$
5.2  Negative values of $n$
Following [7], we chose $n<0$ coprime to $m$. For $n<0$, the differentials $\Omega_{i}^{(j)}(a)$ (6) are not singular at the points $\infty_{1},\dots,\infty_{s}$ at infinity and have poles at the finite ramification points $P_{a_{1}},\dots,P_{a_{N}}$ of the coverings $\zeta:X_{a}\to\mathbb{C}P^{1}.$ As can be seen from definition (6) of the differentials, their behaviour at the pole $P=(\zeta,w)\sim P_{a_{\nu}}$ in terms of the standard local parameter $\xi_{a_{\nu}}$ (2) is
(37) $\Omega_{i}^{(j)}=\frac{w^{-j|n|}}{\zeta-a_{i}}\,d\zeta=\frac{m}{\xi_{a_{\nu}}^{j|n|-m+1}(\xi_{a_{\nu}}^{m}+a_{\nu}-a_{i})}\prod_{h=1,h\neq\nu}^{N}(\xi_{a_{\nu}}^{m}+a_{\nu}-a_{h})^{-j|n|/m}\,d\xi_{a_{\nu}}.$
Thus, $\Omega_{i}^{(j)}$ has a pole of order $j|n|-m+1$ at $P_{a_{\nu}}$ with $\nu\neq i$ if $j|n|\geqslant m$ and of order $j|n|+1$ at $P_{a_{i}}.$ Note that, given that $n$ is coprime to $m$, the residues are nonzero only if $j$ is a multiple of $m$. Let us choose contours of integration in (8) in the form
(38) $\gamma_{l-k}=\sum_{\nu=1}^{N}\frac{c^{\nu}_{l-k}}{2\pi\bf{\rm i}}\gamma_{a_{\nu}}$
where $c^{\nu}_{l-k}\in\mathbb{C}$ are arbitrary constants and where $\gamma_{a_{\nu}}$ is a small contour encircling the point $P_{a_{\nu}}$. Then we obtain from Theorem 1 the following rational solutions to the Schlesinger system given in terms of the residues of $\Omega_{i}^{(j)}(a)$ at $P_{a_{\nu}}$ for some $\theta_{i}\in\mathbb{C}$ with $i=1,\dots,N$ (see Theorem 4 of [7]):
(39) $B^{(i)}_{kl}=\begin{dcases*}\sum_{\nu=1}^{N}c^{\nu}_{l-k}\;\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{i}^{(l-k)}(a),\quad\;\,\text{if }(l-k)/m\in\mathbb{Z},\;\text{ and }\;l>k\\ 0,\qquad\qquad\qquad\qquad\qquad\;\text{if }(l-k)/m\notin\mathbb{Z}\;\text{ or}\;l<k,\\ \theta_{i}-\frac{(k-1)n}{m},\qquad\quad\qquad\,\text{if }\;l=k.\end{dcases*}$
Note, however, that $\sum_{\nu=1}^{N}\gamma_{a_{\nu}}$ corresponds to the sum of residues of the differentials $\Omega_{i}^{j}(a)$ and thus choosing all coefficients $c_{l-k}^{\nu}$ equal in (38) leads to a trivial solution. The nonzero residues from (31), for $(l-k)/m\in\mathbb{Z}$, are explicitly given by [7]
(40) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=m\sum_{k_{1}+\ldots+k_{N}=(l-k)|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-a_{i})^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
for $i=1,\ldots,N,\;i\neq\nu$ and
(41) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{\nu}^{(l-k)}(a)=m\sum_{k_{1}+\ldots+\hat{k}_{\nu}+\ldots+k_{N}=(l-k)|n|/m}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
where the hat symbol over $k_{\nu}$ indicates that the summation index $k_{\nu}$ is absent in the sum in (41). Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (39) are rational functions of $z$ given by the next corollary.
Corollary 10.
Let $n<0,$ and $p,m,N>$ be integers such that ${\rm gcd}(n,m)=1$. Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$. With $c_{j}\in\mathbb{C}$ as in (39), let $S_{j}=S_{j}(z)$ be the rational funtion
(42) $S_{j}(z)=\begin{dcases*}\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\sum_{k_{1}+\ldots+k_{N}=j|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-z)^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-j|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+j|n|/m}},\quad\text{if }j/m\in\mathbb{Z}\\ \hskip 8.5359pt0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\text{if }j/m\notin\mathbb{Z}.\end{dcases*}$
Then the following upper triangular matrix $\Phi(z)$ solves the Fuchsian system (1) with coefficients $B^{(i)},\;i=1,\dots,N,$ being the upper triangular $p\times p$ matrices given by (39):
(43) $\Phi=MD,$
where
(44) $D=\mathrm{diag}\left(\prod_{i=1}^{N}(z-a_{i})^{\beta_{1}^{(i)}},\ldots,\prod_{i=1}^{N}(z-a_{i})^{\beta_{p}^{(i)}}\right),$
and
(45) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{S_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
where, as before, $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
Proof.   Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$. For $n<0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $P_{a_{\nu}}$ for $\nu=1,\dots,N$ as shown at the beginning of Section 5.2. Similarly to the proof of Corollary 8, we need to prove that $\kappa_{j}$ from Definition 2 for the choice of the integration contour $\gamma_{j}$ as in (38) with $l-k=j$ coincides with $S_{j}$ (42). We make an additional assumption that none of the contours $\gamma_{a_{\nu}}$ from (38) contains in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$. With this choice, in terms of the standard local parameter $\xi_{a_{\nu}}$ (2) at $P_{a_{\nu}}$, we have
$\kappa_{j}=-\frac{m}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\,\underset{P_{a_{\nu}}}{\rm res}\,\frac{w^{jn}\mathrm{d}\zeta}{\zeta-z}=\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\;\underset{\xi_{a_{\nu}}=0}{\rm res}\;\frac{\prod_{h=1,h\neq\nu}^{N}(\xi_{a_{\nu}}^{m}+a_{\nu}-a_{h})^{-j|n|/m}\,d\xi_{a_{\nu}}}{\xi_{a_{\nu}}^{j|n|-m+1}(\xi_{a_{\nu}}^{m}+a_{\nu}-z)},$
where the choice of the $m$th root should be made in the same way as in (37). Expanding every factor in Taylor series at $\xi_{a_{\nu}}\sim 0$, we have
$\kappa_{j}=\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\underset{\xi_{a_{\nu}}=0}{\rm res}\;\frac{\Bigl{(}1+\frac{\xi_{a_{\nu}}^{m}}{a_{\nu}-z}\Bigr{)}^{-1}}{\xi_{a_{\nu}}^{j|n|-m+1}(a_{\nu}-z)}\prod_{h=1,h\neq\nu}^{N}\frac{d\xi_{a_{\nu}}}{(a_{\nu}-a_{h})^{j|n|/m}}\Bigl{(}1+\frac{\xi_{a_{\nu}}^{m}}{a_{\nu}-a_{h}}\Bigr{)}^{-j|n|/m}\\ =\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\underset{\xi_{a_{\nu}}=0}{\rm res}\;\frac{d\xi_{a_{\nu}}}{\xi_{a_{\nu}}^{j|n|-m+1}}\sum_{r=0}^{\infty}\Bigl{[}\sum_{k_{1}+\ldots+k_{N}=r}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-z)^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-j|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+j|n|/m}}\Bigr{]}\,\xi_{a_{\nu}}^{rm}.$
Now, to obtain the residue, for each $\nu$, we take the term of the series in $\xi_{a_{\nu}}$ with $r$ such that $rm=j|n|-m$ if such an integer $r$ exists. If $j|n|-m$ is not an integer multiple of $m$, which is equivalent to $j$ not being an integer multiple of $m$ due to the choice of $n$ coprime with $m$, the residue vanishes. This gives us $S_{j}$ from the statement of the corollary. $\Box$
Example 11.
Let $p=3,\;m=N=2$, and $n=-1$. We thus consider the family of curves given by equation $w^{2}=(\zeta-a_{1})(\zeta-a_{2})$ where the associated Riemann surfaces are of genus zero. We define both integration contours to be a constant multiple of the small contour $\gamma_{a_{2}}$ around the point $P_{a_{2}}=(\zeta=a_{2},w=0)$ oriented counterclockwise on the surface: $\gamma_{1}=\gamma_{2}=\frac{c}{2\pi\bf{\rm i}}\gamma_{a_{2}}.$ Note that for every given $z\in\mathbb{C}$ the contour $\gamma_{a_{2}}$ is assumed not to encircle the points $P$ of the surface for which $\zeta(P)=z.$ For any $\theta_{1},\theta_{2},c\in\mathbb{C}$, Corollary 10 gives that the matrix
$\Phi=\left(\begin{array}[]{ccc}(z-a_{1})^{\theta_{1}}(z-a_{2})^{\theta_{2}}&0&\frac{2c}{a_{1}-a_{2}}(z-a_{1})^{\theta_{1}+1}(z-a_{2})^{\theta_{2}}\\ 0&(z-a_{1})^{\theta_{1}+\frac{1}{2}}(z-a_{2})^{\theta_{2}+\frac{1}{2}}&0\\ 0&0&(z-a_{1})^{\theta_{1}+1}(z-a_{2})^{\theta_{2}+1}\end{array}\right)$
satisfies the system
$\frac{d\Phi}{dz}=\left(\begin{array}[]{ccc}\frac{\theta_{1}}{z-a_{1}}+\frac{\theta_{2}}{z-a_{2}}&0&\frac{2c}{(a_{1}-a_{2})^{2}}\left(\frac{1}{z-a_{1}}-\frac{1}{z-a_{2}}\right)\\ 0&\frac{\theta_{1}+\frac{1}{2}}{z-a_{1}}+\frac{\theta_{2}+\frac{1}{2}}{z-a_{2}}&0\\ 0&0&\frac{\theta_{1}+1}{z-a_{1}}+\frac{\theta_{2}+1}{z-a_{2}}\end{array}\right)\Phi.$
Acknowledgments. BP would like to thank Maxime Fortier Bourque (Université de Montréal) for insightful discussions regarding sums over partitions of integers. BP also acknowledges that this research was partially funded by the Fonds de recherche du Québec - section Nature et Technologies (FRQ). VS gratefully acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery grant and from the University of Sherbrooke.

### v2

@@
 The inverse problem, that of reconstructing a function $\Phi(z)$ from its monodromies, is referred to as a Riemann-Hilbert or inverse monodromy problem.
-Alternatively, it may be reformulated as a problem of finding two matrix functions, each of which is holomorphic either inside or outside of a closed contour passing trough the points $a_{1},\dots,a_{N}$ in the complex plane, such that the two functions are related by the multiplication with a “jump” matrix on the contour.
+Alternatively, it may be reformulated as a problem of finding two matrix functions, each of which is holomorphic either inside or outside of a closed contour passing through the points $a_{1},\dots,a_{N}$ in the complex plane, such that the two functions are related by the multiplication with a “jump” matrix on the contour.
 The monodromy matrices associated with the loops around $a_{i}$ may be combined to obtain this jump matrix [4, 17].
@@
 2. Schlesinger system and superelliptic curves
-Let $B^{(1)},B^{(2)},\ldots,B^{(N)}$ be $p\times p$ matrices depending on some complex parameters $a_{1},\ldots,a_{N}$ defined on a ball $D\in\mathbb{C}^{N}\setminus\{(a_{1},\ldots,a_{N})\,\,|\,\,a_{i}=a_{j}\,\text{for some }\,i\neq j\}$.
+Let $B^{(1)},B^{(2)},\ldots,B^{(N)}$ be $p\times p$ matrices depending on some complex parameters $a_{1},\ldots,a_{N}$ defined on a ball $D\subset\mathbb{C}^{N}\setminus\{(a_{1},\ldots,a_{N})\,\,|\,\,a_{i}=a_{j}\,\text{for some }\,i\neq j\}$.
 The Schlesinger system for matrices $B^{(j)}$ has the form
@@
 (4) $\Gamma_{a}=\{(\zeta,w)\in\mathbb{C}^{2}\,|\,w^{m}=\prod_{i=1}^{N}(\zeta-a_{i})\}$
+and we consider their family parametrized by $a=(a_{1},\ldots,a_{N})\in D$ with some fixed positive integers $m$ and $N$.
+The associated compact Riemann surfaces, denoted by $X_{a}$, are obtained by projectivizing and desingularizing the curves (4), see [18] and see also [8, 23] for application to superelliptic curves.
+Note that the algebraic curves (4) are nonsingular.
+The associated projective curves $\hat{\Gamma}_{a}\subset\mathbb{C}P^{2}$ may be singular at the points at infinity denoted collectively by $\{\infty\}$.
+By the theorem of resolution of singularities ([18], Chapter 7) there exists a holomorphic map $\pi:X_{a}\rightarrow\mathbb{C}P^{2}$ with $\pi(X_{a})=\hat{\Gamma}_{a}$ and such that the map
 […]
-$\pi:X_{a}\,\backslash\,\pi^{-1}(\{\infty\})\rightarrow\hat{\Gamma}_{a}\,\backslash\,\{\infty\}$
 is biholomorphic.
@@
 (5) $\xi_{\infty_{j}}(P)=(\zeta(P))^{-\frac{1}{m_{1}}},$ $\mbox{if}\quad P\sim\infty_{j},\quad\mbox{ where }m=sm_{1}.$
-$\xi_{Q}(P)=\zeta(P)-\zeta(Q),$ $\mbox{if}\quad P\sim Q\quad\mbox{and $Q$ is a regular point.}$.
+$\xi_{Q}(P)=\zeta(P)-\zeta(Q),$ $\mbox{if}\quad P\sim Q\quad\mbox{and $Q$ is a regular point.}$
 We thus consider the family of compact superelliptic Riemann surfaces $X_{a}$, $a\in D,$ associated with the family of algebraic curves (4).
@@
 Let $\Omega_{i}^{j}(a)$ be the meromorphic differential defined by (6) on $X_{a}$.
-Then a set of upper triangular $p\times p$ matrices $B^{(i)}$, $i=1,...,N$ defined by the following three conditions provides a solution to the Schlesinger system (3):
+Then a set of upper triangular $p\times p$ matrices $B^{(i)}$, i=1,,Ni=1,...,N defined by the following three conditions provides a solution to the Schlesinger system (3):
   • $B^{(i)}_{kl}=0$ if $k>l;$
-  • The eigenvalues of each matrix $B^{(i)}$, $i=1,...,N$, satisfy
+  • The eigenvalues of each matrix $B^{(i)}$, i=1,,Ni=1,...,N, satisfy
     (7) $B^{(i)}_{kk}-B^{(i)}_{k+1,k+1}=\frac{n}{m},\quad\text{with }k=1,\ldots,p-1;$
@@
     1. (a)
-      $H_{1}(X_{a},\mathbb{C})$ if $m$, $N$ are coprime and $n>0$,
+      $H_{1}(X_{a},\mathbb{C})$ if $m$, NN are coprime and $n>0$,
     2. (b)
-      $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s}\},\mathbb{C})$ if $m$, $N$ are not coprime, $s={\rm gcd}(m,N)$, and $n>0$.
+      $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s}\},\mathbb{C})$ if $m$, NN are not coprime, $s={\rm gcd}(m,N)$, and $n>0$.
     3. (c)
-      $H_{1}(X_{a}\setminus\{P_{a_{1}},\ldots,P_{a_{N}}\},{\mathbb{C}})$ if $n<0\,.$
+      $H_{1}(X_{a}\setminus\{P_{a_{1}},\ldots,P_{a_{N}}\},{\mathbb{C}})$ if n<0.n<0\,.
 Note that for some contours, the integrals in (8) vanish giving possibly trivial solutions to the Schlesinger system.
@@
 Theorem 4.
-Let $p,\;n,\;m,\;N$ and $X_{a},a\in D$ be as in Theorem 1.
-Let $z\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ and let $\gamma_{j}$ be a closed contour on $X_{a}$ from Theorem 1 not passing through any point $P\in X_{a}$ for which $\zeta(P)=z$.
+Let p,n,m,Np,\;n,\;m,\;N and $X_{a},a\in D$ be as in Theorem 1.
+Let $z_{0}\in\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ and let $\gamma_{1},\dots,\gamma_{p-1}$ be closed contours on $X_{a}$ from Theorem 1 not passing through any point $P\in X_{a}$ for which $\zeta(P)=z_{0}$.
 More precisely, let $\gamma_{j}$ from Theorem 1 be seen as an element of
   1. (a)
-    $H_{1}(X_{a}\setminus\{P\in X_{a}\,|\,\zeta(P)=z\},\mathbb{C})$ if $m$, $N$ are coprime and $n>0$,
-  2. (a)
-    $H_{1}(X_{a}\setminus(\{\infty_{1},\ldots,\infty_{s}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z\}),\mathbb{C})$ if $n>0$ and $s={\rm gcd}(m,N)>1$,
+    $H_{1}(X_{a}\setminus\{P\in X_{a}\,|\,\zeta(P)=z_{0}\},\mathbb{C})$ if $m$, NN are coprime and $n>0$,
   2. (b)
-    $H_{1}(X_{a}\setminus(\{P_{a_{1}},\ldots,P_{a_{N}}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z\}),{\mathbb{C}})$ if $n<0$,
+    $H_{1}(X_{a}\setminus(\{\infty_{1},\ldots,\infty_{s}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z_{0}\}),\mathbb{C})$ if $n>0$ and $s={\rm gcd}(m,N)>1$,
+  3. (c)
+    $H_{1}(X_{a}\setminus(\{P_{a_{1}},\ldots,P_{a_{N}}\}\cup\{P\in X_{a}\,|\,\zeta(P)=z_{0}\}),{\mathbb{C}})$ if $n<0$,
 independent of small variations of $a\in D$.
-For such a choice of the contour $\gamma_{j}$, let $\kappa_{j}$ be a function of $z$ given by Definition 2.
+Let $\mathcal{Z}\subset\mathbb{C}\setminus\{a_{1},\dots,a_{N}\}$ be a small neighbourhood of $z_{0}$ not intersecting projections on the $\zeta$-sphere of the contours $\gamma_{1},\dots,\gamma_{p-1}$.
+For such a choice of the contour $\gamma_{j}$, let $\kappa_{j}$ be a function of $z\in\mathcal{Z}$ given by Definition 2.
 Let $B^{(i)},\;i=1,\dots,N,$ be upper triangular $p\times p$ matrices given by Theorem 1 with eigenvalues
 $\beta_{k}^{(i)}=B^{(i)}_{kk}=\theta_{i}-\frac{(k-1)n}{m},\quad k\in\{1,\ldots,p\},$
-where $\theta_{i}\in\mathbb{R}$. Then, a fundamental upper triangular matrix solution $\Phi(z)$ to the Fuchsian system (1) with coefficients being the matrices $B^{(i)}$, $i=1,\ldots,N$, is given by
+where $\theta_{i}\in\mathbb{R}$.
+Then, for $z\in\mathcal{Z}$ a fundamental upper triangular matrix solution $\Phi(z)$ to the Fuchsian system (1) with coefficients being the matrices $B^{(i)}$, $i=1,\ldots,N$, is given by
 (12) $\Phi=MD,$
@@
 and
-(14) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{\kappa_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
-where $q\vdash(l-k)$ represents the partitions of the integer $l-k$, and $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
+[…]
+where $q\vdash(l-k)$ represents the partitions of the integer lkl-k, and $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
 Note that as $z$ varies, the contours $\gamma_{j}$ may be adjusted to avoid the points $P\in X_{a}$ with $\zeta(P)=z$ while staying in the same homology class from Theorem 4.
@@
 However, each set of contours $\gamma_{1},\dots,\gamma_{p-1}$ corresponding to a fixed solution $B^{(1)},\dots,B^{(N)}$ (7), (8) of the Schlesinger system and thus to a fixed Fuchsian system (1), gives rise to several different solutions of the Fuchsian system (corresponding to different initial conditions for (1)).
-This happens because, as stated in Theorem 4, the contours $\gamma_{j}$ are elements of the homology of the surface punctured at the points $P\in X_{a}$ for which $\zeta(P)=z.$ In other words, a given contour $\gamma_{j}$ in $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s},P_{a_{1}},\ldots,P_{a_{N}}\},\mathbb{C})$ will give different values of the integral $\kappa_{j}$ (9) depending on the positions of the points of the surface satisfying $\zeta(P)=z_{0}$ relative to the contour $\gamma_{j}.$ Here we assume that initial value for $\Phi$ is given at $z=z_{0}$.
+This happens because, as stated in Theorem 4, the contours $\gamma_{j}$ are elements of the homology of the surface punctured at the points $P\in X_{a}$ for which $\zeta(P)=z_{0}.$ In other words, a given contour $\gamma_{j}$ in $H_{1}(X_{a}\setminus\{\infty_{1},\ldots,\infty_{s},P_{a_{1}},\ldots,P_{a_{N}}\},\mathbb{C})$ will give different values of the integral $\kappa_{j}$ (9) depending on the positions of the points of the surface satisfying $\zeta(P)=z_{0}$ relative to the contour $\gamma_{j}.$ Here we assume that initial value for $\Phi$ is given at $z=z_{0}$.
 Let us illustrate this with a simple example.
@@
     The contour $\gamma$ contains $t<m$ points $P_{z}^{(1)},\dots,P_{z}^{(t)}$ in its interior. In this case, we have for $\kappa_{j}$
-    (19) $\kappa_{j}:=-\frac{m}{jn}\oint_{\gamma}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\underset{P_{z}^{(r)}}{\rm res}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\epsilon_{r}\Pi^{\frac{jn}{m}},$
+    (19) $\kappa_{j}:=-\frac{m}{jn}\oint_{\gamma}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi{\mathrm{i}}\sum_{r=1}^{t}\underset{P_{z}^{(r)}}{\rm res}\frac{w^{jn}}{\zeta-z}\mathrm{d}\zeta=-\frac{m}{jn}2\pi{\mathrm{i}}\sum_{r=1}^{t}\epsilon_{r}\Pi^{\frac{jn}{m}},$
     where $\epsilon_{r}$ are different $m$th roots of unity.
-Let us, for simplicity of notation, denote $\alpha_{j}:=-\frac{m}{jn}2\pi\bf{\rm i}\sum_{r=1}^{t}\epsilon_{r}$ so that $\kappa_{j}=\alpha_{j}\Pi^{\frac{jn}{m}}.$ Now, for the entries above the diagonal of the matrix $M$ given by (14), we have:
+Let us, for simplicity of notation, denote $\alpha_{j}:=-\frac{m}{jn}2\pi{\mathrm{i}}\sum_{r=1}^{t}\epsilon_{r}$ so that $\kappa_{j}=\alpha_{j}\Pi^{\frac{jn}{m}}.$ Now, for the entries above the diagonal of the matrix $M$ given by (14), we have:
     $M_{kl}=\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{\left(\alpha_{j}\Pi^{\frac{jn}{m}}\right)^{\sigma_{q}(j)}}{\sigma_{q}(j)!}=\hat{\alpha}_{l-k}\Pi^{\frac{(l-k)n}{m}}$
@@
   1. (1)
-    $\frac{\mathrm{d}}{\mathrm{d}z}D=D_{z}=\mathrm{diag}\left(\sum_{i=1}^{N}\frac{\beta_{1}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{1}^{(i)}},\ldots,\sum_{i=1}^{N}\frac{\beta_{p}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{p}^{(i)}}\right)$,
+    $\frac{\mathrm{d}}{\mathrm{d}z}D=D_{z}=\mathrm{diag}\left(\sum_{i=1}^{N}\frac{\beta_{1}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{1}^{(j)}},\ldots,\sum_{i=1}^{N}\frac{\beta_{p}^{(i)}}{z-a_{i}}\prod_{j=1}^{N}(z-a_{j})^{\beta_{p}^{(j)}}\right)$,
   2. (2)
     $\frac{\mathrm{d}}{\mathrm{d}z}M=M_{z}$ for which the $(kl)$-entry is
-    $(M_{z})_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\sum_{r=1}^{l-k}\left(\frac{\kappa_{r}^{\sigma_{q}(r)-1}}{(\sigma_{q}(r)-1)!}\kappa_{r}^{\prime}\prod_{j=1,j\neq r}^{l-k}\frac{\kappa_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\right)\quad&\text{when } $l>k$,\\ \quad 0\quad&\text{when } $l\leq k$,\end{dcases*}$
+[…]
   3. (3)
-    $D^{-1}=\mathrm{diag}\left(\prod_{j=1}^{N}(z-a_{j})^{-\beta_{1}^{(i)}},\ldots,\prod_{j=1}^{N}(z-a_{j})^{-\beta_{p}^{(i)}}\right)$, and
+    $D^{-1}=\mathrm{diag}\left(\prod_{j=1}^{N}(z-a_{j})^{-\beta_{1}^{(j)}},\ldots,\prod_{j=1}^{N}(z-a_{j})^{-\beta_{p}^{(j)}}\right)$, and
   4. (4)
     the inverse of the matrix $M$ has the following entries
-    (21) $(M^{-1})_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{(-\kappa_{j})^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$.\end{dcases*}$
+[…]
 Proof.   Items (1) - (3) of the lemma being obtained by straightforward observation, let us prove item (4). We want to show that
@@
 $\sum_{K+L=T}\left(\sum_{r_{1}\vdash K}\prod_{j=1}^{K}\frac{\kappa_{j}^{\sigma_{r_{1}}(j)}}{\sigma_{r_{1}}(j)!}\right)\left(\sum_{r_{2}\vdash L}\prod_{j=1}^{L}\frac{(-\kappa_{j})^{\sigma_{r_{2}}(j)}}{\sigma_{r_{2}}(j)!}\right)\\ =\sum_{q\vdash T}\prod_{j=1}^{T}\frac{1}{\sigma_{q}(j)!}\sum_{\sigma_{r_{1}}+\sigma_{r_{2}}=\sigma_{q}(j)}\frac{\sigma_{q}(j)!}{\sigma_{r_{1}}!\sigma_{r_{2}}!}\kappa_{j}^{\sigma_{r_{1}}}(-\kappa_{j})^{\sigma_{r_{2}}}=\sum_{q\vdash T}\prod_{j=1}^{T}\frac{1}{\sigma_{q}(j)!}(\kappa_{j}-\kappa_{j})^{\sigma_{q}(j)}=0.$
-For $T=l-k$, this implies (22) with $l>k$ and thus proves that (21) defines the inverse of the matrix $M$ (14). $\Box$ We now have all the necessary tools to prove Theorem 4.
+For T=lkT=l-k, this implies (22) with $l>k$ and thus proves that (21) defines the inverse of the matrix $M$ (14). $\Box$ We now have all the necessary tools to prove Theorem 4.
 Proof of Theorem 4. We want to show that $\Phi=MD$ where both $M$ and $D$ are defined in Theorem 4, is a fundamental solution of the Fuchsian system (1), i.e.
@@
 Let us consider the entries of $M_{z}M^{-1}$ lying above the main diagonal.
-Fix $k<l$ and set $j=l-k$.
+Fix $k<l$ and set j=lkj=l-k.
 By Lemma 7, for $r\geqslant 1$, we have
@@
 Let us now consider the contribution, denoted by $C(q,t),$ to (26) of a fixed partition $q\vdash j$ and of a fixed value of $t$.
-Introduce the vector $\sigma_{q}=(\sigma_{q}(1),\dots,\sigma(j))$, which defines the partition $q$ uniquely.
+Introduce the vector $\sigma_{q}=(\sigma_{q}(1),\dots,\sigma_{q}(j))$, which defines the partition $q$ uniquely.
 We can rewrite the contribution of $q$ and $t$ in (26) as a sum over all decompositions of $\sigma_{q}$ into the sum of two vectors.
-Then, similarly to (23), this becomes the product of sums (we write $\sigma_{k}(i)$ for the $i$th component of the vector $\sigma_{k}$):
-$C(q,t)=\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\end{subarray}}\sigma_{q_{1}}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{q_{2}}(i)}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}=\sum_{\sigma_{1}+\sigma_{2}=\sigma_{q}}\sigma_{1}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{2}(i)}\binom{\sigma_{q}(i)}{\sigma_{1}(i)}$
+Then, similarly to (23), this becomes the product of sums (we write $v_{k}(i)$ for the $i$th component of the vector $v_{k}$):
+$C(q,t)=\sum_{\begin{subarray}{c}(q_{1},q_{2})\\ q_{1}+q_{2}=q\end{subarray}}\sigma_{q_{1}}(t)\;\prod_{i=1}^{j}(-1)^{\sigma_{q_{2}}(i)}\binom{\sigma_{q}(i)}{\sigma_{q_{1}}(i)}=\sum_{v_{1}+v_{2}=\sigma_{q}}v_{1}(t)\;\prod_{i=1}^{j}(-1)^{v_{2}(i)}\binom{\sigma_{q}(i)}{v_{1}(i)}$
 (27) $=\left(\sum_{x_{t}=0}^{\sigma_{q}(t)}x_{t}\binom{\sigma_{q}(t)}{x_{t}}(-1)^{\sigma_{q}(t)-x_{t}}\right)\prod_{\begin{subarray}{c}i=1\\ i\neq t\end{subarray}}^{j}\left(\sum_{x_{i}=0}^{\sigma_{q}(i)}\binom{\sigma_{q}(i)}{x_{i}}(-1)^{\sigma_{q}(i)-x_{i}}\right).$
-If $\sigma_{q}(i)\neq 0$, the sum over $x_{i}$ for $i\neq t$ equals $(1+x)^{\sigma_{q}(i)}$ evaluated at $x=-1$ and thus vanishes.
+If $\sigma_{q}(i)\neq 0$, the sum over $x_{i}$ for $i\neq t$ equals $(1+x)^{\sigma_{q}(i)}$ evaluated at x=1x=-1 and thus vanishes.
 If $\sigma_{q}(i)=0$, the sum over $x_{i}$ is equal to $1$.
-The sum over $x_{t}$, if $\sigma_{q}(t)>1$, can be rewritten as the derivative of $(1+x)^{\sigma_{q}(t)}$ evaluated at $-1$, therefore it is equal to $1$ if $\sigma_{q}(t)=1$ and 00 otherwise.
+The sum over $x_{t}$, if $\sigma_{q}(t)>0$, can be rewritten as the derivative of $(1+x)^{\sigma_{q}(t)}$ evaluated at 1-1, therefore it is equal to $1$ if $\sigma_{q}(t)=1$ and 0 otherwise.
 Thus, $C(q,t)$ is non-zero if and only if $q=j$ and $t=j$, that is if $q$ is the partition of $j$ containing one part only, in which case $C(q,t)=1$.
@@
 $(MD_{z}D^{-1}M^{-1})_{kl}$ $=\sum_{i=1}^{N}\sum_{h=k}^{l}M_{kh}\,\frac{\beta_{h}^{(i)}}{\;z-a_{i}\!}\,(M^{-1})_{hl}$
-$=\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}}\!\sum_{h=k}^{l}M_{kh}(M^{-1})_{hl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{h=k}^{l}(h-k)M_{kh}(M^{-1})_{hl}$
-$=\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}}\delta_{kl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{r=0}^{l-k}rM_{k,k+r}(M^{-1})_{k+r,l}.$
-Comparing with our above derivation, we see that the second term can be rewritten using partitions similarly to (26), where $j=l-k$.
-The factor of $r$ corresponds to the sum of all the parts of the partition $q_{1}$.
+$=\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}}\!\sum_{h=k}^{l}M_{kh}(M^{-1})_{hl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{h=k}^{l}(h-k)M_{kh}(M^{-1})_{hl}$
+$=\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}}\delta_{kl}\;-\;\frac{n}{m}\sum_{i=1}^{N}\frac{1}{z-a_{i}}\sum_{r=0}^{l-k}rM_{k,k+r}(M^{-1})_{k+r,l}.$
+Comparing with our above derivation, we see that the second term can be rewritten using partitions similarly to (26), where j=lkj=l-k.
+The factor of $r$ corresponds to the sum of all the parts of the partition $q_{1}$, that is $r=\sum_{t=1}^{j}t\,\sigma_{q_{1}}(t)$.
 More precisely, using (14) for $M$ as well as formulas from (25) to (3), we obtain
@@
 $=\sum_{q\vdash j}\left(\prod_{i=1}^{j}\frac{\left(\kappa_{i}\right)^{\sigma_{q}(i)}}{\sigma_{q}(i)!}\right)\sum_{t=1}^{j}t\,C(q,t).$
-Given that $C(q,t)$ is non-zero only in the case of one term partition $q=j=l-k$ and $t=j,$ in which case $C(j,j)=1$, we obtain that quantity (28) is equal to $j\kappa_{j}$ and thus
-$(MD_{z}D^{-1}M^{-1})_{kl}=\begin{cases}0,&l<k,\\ \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}},&l=k,\\ \displaystyle-\,\frac{(l-k)n}{m}\,\kappa_{\,l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
+Given that $C(q,t)$ is non-zero only in the case of one term partition q=j=lkq=j=l-k and $t=j,$ in which case $C(j,j)=1$, we obtain that quantity (28) is equal to $j\kappa_{j}$ and thus
+$(MD_{z}D^{-1}M^{-1})_{kl}=\begin{cases}0,&l<k,\\ \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}},&l=k,\\ \displaystyle-\,\frac{(l-k)n}{m}\,\kappa_{\,l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
 Summing the two contributions into (24), we have
-$(\Phi_{z}\Phi^{-1})_{kl}=\begin{cases}0,&l<k,\\[3.0pt] \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(l)}}{z-a_{i}},&l=k,\\[4.0pt] \kappa_{l-k}^{\prime}-\displaystyle\frac{({l-k})n}{m}\,\kappa_{l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
+$(\Phi_{z}\Phi^{-1})_{kl}=\begin{cases}0,&l<k,\\[3.0pt] \displaystyle\sum_{i=1}^{N}\frac{\beta_{k}^{(i)}}{z-a_{i}},&l=k,\\[4.0pt] \kappa_{l-k}^{\prime}-\displaystyle\frac{({l-k})n}{m}\,\kappa_{l-k}\sum_{i=1}^{N}\frac{1}{z-a_{i}},&l>k.\end{cases}$
 Recalling that from Lemma 3, for each $l-k\geq 1$,
@@
 Thus we obtain that
-$\mathcal{M}_{i}=\mathrm{diag}\left(e^{2\pi\mathrm{i}\beta_{1}^{(i)}},\ldots,e^{2\pi\mathrm{i}\beta_{p}^{(i)}}\right)=\mathrm{diag}\left(e^{2\pi\mathrm{i}\theta_{i}},e^{2\pi\mathrm{i}\theta_{i}-\frac{n}{m}},\ldots,e^{2\pi\mathrm{i}\theta_{i}-\frac{(p-1)n}{m}}\right).$
+$\mathcal{M}_{i}=\mathrm{diag}\left(e^{2\pi{\mathrm{i}}\beta_{1}^{(i)}},\ldots,e^{2\pi{\mathrm{i}}\beta_{p}^{(i)}}\right)=\mathrm{diag}\left(e^{2\pi{\mathrm{i}}\theta_{i}},e^{2\pi{\mathrm{i}}(\theta_{i}-\frac{n}{m})},\ldots,e^{2\pi{\mathrm{i}}(\theta_{i}-\frac{(p-1)n}{m})}\right).$
 Case 2.
@@
 We give the monodromy matrices for the case $p=3$:
-$\mathcal{M}_{i}={\rm e}^{2\pi\bf{\rm i}\theta_{i}}\begin{pmatrix}1&\hat{\alpha}_{1}\left(1-{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)&\hat{\alpha}_{2}\left(1-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)-\hat{\alpha}_{1}^{2}\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}&\hat{\alpha}_{1}\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\end{pmatrix}.$
+$\mathcal{M}_{i}={\rm e}^{2\pi{\mathrm{i}}\theta_{i}}\begin{pmatrix}1&\hat{\alpha}_{1}\left(1-{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}\right)&\hat{\alpha}_{2}\left(1-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)-\hat{\alpha}_{1}^{2}\left({\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}&\hat{\alpha}_{1}\left({\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\end{pmatrix}$.
 Case 3.
@@
 The monodromy matrices are for $i=1,\dots,N$
-$\mathcal{M}_{i}={\rm e}^{2\pi\bf{\rm i}\theta_{i}}\begin{pmatrix}1&\alpha\left(1-{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)&\frac{1}{2}\alpha^{2}\left(1+{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}-2{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}&\alpha\left({\rm e}^{-\frac{2\pi\bf{\rm i}n}{m}}-{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi\bf{\rm i}n}{m}}\end{pmatrix}.$
-We see that in all cases the monodromy matrices are independent of small variations of $a_{i}$, $i=1,\dots,N$, and are conjugate to ${\rm exp}\{2\pi\bf{\rm i}\,B^{(i)}\}$ as expected according to the general theory of Fuchsian systems, see for example, [16].
-Note that our solutions also transform when $z$ goes around the point at infinity, and the corresponding monodromy matrix $\mathcal{M}_{\infty}$ is conjugate to ${\rm exp}\{2\pi\bf{\rm i}\,B^{(\infty)}\}$ where $B^{(\infty)}=\sum_{i=1}^{N}B^{(i)}$ is a nonzero diagonal matrix, see the end of Section 2.
+$\mathcal{M}_{i}={\rm e}^{2\pi{\mathrm{i}}\theta_{i}}\begin{pmatrix}1&\alpha\left(1-{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}\right)&\frac{1}{2}\alpha^{2}\left(1+{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}-2{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}\right)\\ 0&{\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}&\alpha\left({\rm e}^{-\frac{2\pi{\mathrm{i}}n}{m}}-{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\right)\\ 0&0&{\rm e}^{-\frac{4\pi{\mathrm{i}}n}{m}}\end{pmatrix}$.
+We see that in all cases the monodromy matrices are independent of small variations of $a_{i}$, $i=1,\dots,N$, and are conjugate to ${\rm exp}\{2\pi{\mathrm{i}}\,B^{(i)}\}$ as expected according to the general theory of non-resonant Fuchsian systems, see Chapter IV, Theorem 4.1 of [6].
+Note that our solutions also transform when $z$ goes around the point at infinity, and the corresponding monodromy matrix $\mathcal{M}_{\infty}$ is conjugate to ${\rm exp}\{2\pi{\mathrm{i}}\,B^{(\infty)}\}$ where $B^{(\infty)}=\sum_{i=1}^{N}B^{(i)}$ is a nonzero diagonal matrix, see the end of Section 2.
 A loop around $z=\infty$ is given by a product of loops $\rho_{i}$ and thus $\mathcal{M}_{\infty}$ is obtained as the corresponding product of the matrices $\mathcal{M}_{i}$.
-5.  Some rational solutions
+5. Some solutions $\Phi=MD$ with rational $M$
 As mentioned after Theorem 1, the contours of integration may be chosen in a way for the solutions of the Schlesinger system from Theorem 1 to be given by residues of the differentials (6).
@@
 For coprime $m$ and $N$, there is only one pole at infinity and thus the residue there is zero.
-For $n<0$, all poles are at the finite ramification points $P_{a_{1}},\dots,P_{a_{N}}.$ In these cases we obtain rational solutions to the Schlesinger system and to the Fuchsian system as well.
+For $n<0$, all poles are at the finite ramification points $P_{a_{1}},\dots,P_{a_{N}}.$ In these cases we obtain, respectively, polynomial and rational solutions to the Schlesinger system.
+For the corresponding solutions $\Phi=MD$ to the Fuchsian system, the matrix $M$ is rational.
+The solution $\Phi$ is algebraic for rational constants $\theta_{i}$, $i=1,\dots,N,$ from Theorem 4.
 5.1  Positive values of $n$
@@
 The behaviour of these differentials at $P=(\zeta,w)\sim\infty_{\alpha}$ for $\alpha=1,\dots,s$ in terms of the local parameter $\xi:=\xi_{\infty_{\alpha}}$ given by (2) is
-(29) $\Omega_{i}^{(j)}(P)=\frac{w^{jn}d\zeta}{\zeta-a_{i}}=\frac{\nu_{jn\alpha}\prod_{i=1}^{N}(1-a_{i}\xi^{m_{1}})^{jn/m}}{\xi^{jnN_{1}+1}(1-a_{i}\xi^{m_{1}})}\,d\xi,\qquad\nu_{jn\alpha}=-m_{1}\,e^{2\pi\bf{\rm i}jn(\alpha-1)/s}.$
+(29) $\Omega_{i}^{(j)}(P)=\frac{w^{jn}d\zeta}{\zeta-a_{i}}=\frac{\nu_{jn\alpha}\prod_{i=1}^{N}(1-a_{i}\xi^{m_{1}})^{jn/m}}{\xi^{jnN_{1}+1}(1-a_{i}\xi^{m_{1}})}\,d\xi,\qquad\nu_{jn\alpha}=-m_{1}\,e^{2\pi{\mathrm{i}}jn(\alpha-1)/s}.$
 We thus have a pole of order $jnN_{1}+1$ at $P=\infty_{\alpha}$ for each $\Omega_{i}^{(j)}$. As can be seen from (29) the residue at the pole is nonzero only if $j$ is a multiple of $m_{1}$.
@@
 Let $\gamma_{\infty_{\alpha}}$ be a small contour encircling $\infty_{\alpha}$.
-Choosing now $\gamma_{l-k}=\frac{c_{l-k}}{2\pi\bf{\rm i}A_{l-k}}\gamma_{\infty_{\alpha}}$ for some $c_{l-k}\in\mathbb{C}$, we have a polynomial solution to the Schlesinger system given, for some $\theta_{i}\in\mathbb{C}$, by (see Theorem 3 of [7])
+Choosing now $\gamma_{l-k}=\frac{c_{l-k}}{2\pi{\mathrm{i}}A_{l-k}}\gamma_{\infty_{\alpha}}$ for some $c_{l-k}\in\mathbb{C}$, we have a polynomial solution to the Schlesinger system given, for some $\theta_{i}\in\mathbb{C}$, by (see Theorem 3 of [8])
 (31) $B^{(i)}_{kl}=\begin{dcases*}\frac{c_{l-k}}{A_{l-k}}\;\underset{\infty_{\alpha}}{\rm res}\,\Omega_{i}^{(l-k)}(a),\quad\text{if }(l-k)/m_{1}\in\mathbb{Z}\;\text{ and }\;l>k\\ 0,\qquad\qquad\qquad\qquad\;\;\,\text{if }(l-k)/m_{1}\notin\mathbb{Z}\;\text{ or }\;l<k\\ \theta_{i}-\frac{(k-1)n}{m},\qquad\quad\;\;\,\text{if }\;l=k.\end{dcases*}$
 The nonzero residue from (31), for $(l-k)/m_{1}\in\mathbb{Z}$, is explicitly given by [8]
-(32) $\underset{\infty_{\alpha}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=A_{l-k}\sum_{k_{1}+\ldots+k_{N}+q=N(l-k)\frac{n}{m}}(-1)^{q}{(l-k)\frac{n}{m}\choose k_{1}}\ldots{(l-k)\frac{n}{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}a_{i}^{q},$
+(32) $\underset{\infty_{\alpha}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=A_{l-k}\sum_{k_{1}+\ldots+k_{N}+t=N(l-k)\frac{n}{m}}(-1)^{t}{(l-k)\frac{n}{m}\choose k_{1}}\ldots{(l-k)\frac{n}{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}a_{i}^{t},$
 with the generalized binomial coefficients
-(${\beta\choose j}=\frac{\beta(\beta-1)\cdots(\beta-j+1)}{j!},\qquad{\beta\choose 0}=1$
+${\beta\choose j}=\frac{\beta(\beta-1)\cdots(\beta-j+1)}{j!},\qquad{\beta\choose 0}=1$
 defined for $\beta\in\mathbb{R}$ and $j\in\mathbb{N}$.
-Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (31) are rational functions of $z$ given by the next corollary.
+Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (31) are functions of $z$ given by the next corollary.
+These solutions are algebraic for rational values of $\theta_{i}$.
 Corollary 8.
 Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and $p,n,m,N$ be strictly positive integers such that ${\rm gcd}(n,m)=1$ and $s={\rm gcd}(m,N)>1$.
-With $c_{j}\in\mathbb{C}$ as in (31), let $R_{j}=R_{j}(z)$ be the rational funtion
-(33) $R_{j}(z)=\begin{dcases*}-\frac{m\,c_{j}}{jn}\sum_{k_{1}+\ldots+k_{N}+q=N\frac{jn}{m}}(-1)^{q}{{jn}/{m}\choose k_{1}}\ldots{{jn}/{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{q},\quad\text{if }j/m_{1}\in\mathbb{Z}\\ 0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\;\text{if }j/m_{1}\notin\mathbb{Z}.\end{dcases*}$
+With $c_{j}\in\mathbb{C}$ as in (31), let $R_{j}=R_{j}(z)$ be the polynomial
+(33) $R_{j}(z)=\begin{dcases*}-\frac{m\,c_{j}}{jn}\sum_{k_{1}+\ldots+k_{N}+t=N\frac{jn}{m}}(-1)^{t}{{jn}/{m}\choose k_{1}}\ldots{{jn}/{m}\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{t},\quad\text{if }j/m_{1}\in\mathbb{Z}\\ 0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\;\text{if }j/m_{1}\notin\mathbb{Z}.\end{dcases*}$
 Denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$ and $i=1,\dots,N$.
-Then the following upper triangular matrix $\Phi(z)$ solves the Fuchsian system (1) with coefficients $B^{(i)}$ being the upper triangular $p\times p$ matrices given by (31).
+Then the following upper triangular matrix $\Phi(z)$ solves the Fuchsian system (1) with coefficients $B^{(i)}$ being the upper triangular $p\times p$ matrices given by (31):
 (34) $\Phi=MD,$
@@
 and
-(36) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{R_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
+[…]
 where $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
@@
 For $n>0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $\infty_{1},\dots,\infty_{s}$ as explained at the beginning of Section 5.1.
-Comparing the statements of Corollary 8 and Theorem 4, we only need to prove that $\kappa_{j}$ from Definition 2 coincides with $R_{j}$ (33) for the choice of the integration contour $\gamma_{j}=\frac{c_{j}}{2\pi\bf{\rm i}A_{j}}\gamma_{\infty_{\alpha}}$ with $\gamma_{\infty_{\alpha}}$ being a small contour encircling $\infty_{\alpha}\in X_{a}$ and not containing in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$.
+Comparing the statements of Corollary 8 and Theorem 4, we only need to prove that $\kappa_{j}$ from Definition 2 coincides with $R_{j}$ (33) for the choice of the integration contour $\gamma_{j}=\frac{c_{j}}{2\pi{\mathrm{i}}A_{j}}\gamma_{\infty_{\alpha}}$ with $\gamma_{\infty_{\alpha}}$ being a small contour encircling $\infty_{\alpha}\in X_{a}$ and not containing in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$.
 With this choice and $m_{1},N_{1}$ as above, we have
@@
 where $\xi:=\xi_{\infty_{\alpha}}$ is the local parameter at $\infty_{\alpha}$ given by (2) and $\nu_{jn\alpha}$ as in (29). Therefore, expanding every factor in Taylor series in $\xi\sim 0$,
-$\kappa_{j}$ $=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{k_{1}=0}^{\infty}{jn/m\choose k_{1}}(-a_{1}\xi^{m_{1}})^{k_{1}}\ldots\sum_{k_{N}=0}^{\infty}{jn/m\choose k_{N}}(-a_{N}t^{m_{1}})^{k_{N}}\sum_{q=0}^{\infty}(z\xi^{m_{1}})^{q}$
-$=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{r=0}^{\infty}\Bigl{[}\sum_{k_{1}+\ldots+k_{N}+q=r}(-1)^{r-q}{jn/m\choose k_{1}}\ldots{jn/m\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{q}\Bigr{]}\,\xi^{rm_{1}}.$
+$\kappa_{j}$ $=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{k_{1}=0}^{\infty}{jn/m\choose k_{1}}(-a_{1}\xi^{m_{1}})^{k_{1}}\ldots\sum_{k_{N}=0}^{\infty}{jn/m\choose k_{N}}(-a_{N}\xi^{m_{1}})^{k_{N}}\sum_{t=0}^{\infty}(z\xi^{m_{1}})^{t}$
+$=$ $-\frac{m\,c_{j}}{jnA_{j}}\,\underset{\xi=0}{\rm res}\,\frac{\nu_{jn\alpha}d\xi}{\xi^{jnN_{1}+1}}\sum_{r=0}^{\infty}\Bigl{[}\sum_{k_{1}+\ldots+k_{N}+t=r}(-1)^{r-t}{jn/m\choose k_{1}}\ldots{jn/m\choose k_{N}}a_{1}^{k_{1}}\ldots a_{N}^{k_{N}}z^{t}\Bigr{]}\,\xi^{rm_{1}}.$
 To obtain the residue we need to take the term of the series corresponding to the value of $r$ such that $jnN_{1}=rm_{1}$. Note that the constant $A_{j}$ cancels out and we obtain $R_{j}$ (33). $\Box$
 Example 9.
-Let us set, for simplicity, $\theta_{1}=\theta_{2}=\frac{1}{4}$, $p=2$, and $n=1,\;m=2,\;N=2$ so that $s=2$.
+Let us set, for simplicity, $\theta_{1}=\theta_{2}=\frac{1}{4}$, $p=2$, and n=1,m=2,N=2n=1,\;m=2,\;N=2 so that $s=2$.
 In this case, $\beta_{1}=\frac{1}{4}$ and $\beta_{2}=-\frac{1}{4}$ and the polynomial in $a_{1},\,a_{2}$ solution (31) to the Schlesinger system is
@@
 (37) $\Omega_{i}^{(j)}=\frac{w^{-j|n|}}{\zeta-a_{i}}\,d\zeta=\frac{m}{\xi_{a_{\nu}}^{j|n|-m+1}(\xi_{a_{\nu}}^{m}+a_{\nu}-a_{i})}\prod_{h=1,h\neq\nu}^{N}(\xi_{a_{\nu}}^{m}+a_{\nu}-a_{h})^{-j|n|/m}\,d\xi_{a_{\nu}}.$
-Thus, $\Omega_{i}^{(j)}$ has a pole of order $j|n|-m+1$ at $P_{a_{\nu}}$ with $\nu\neq i$ if $j|n|\geqslant m$ and of order $j|n|+1$ at $P_{a_{i}}.$ Note that, given that $n$ is coprime to $m$, the residues are nonzero only if $j$ is a multiple of $m$.
+Thus, $\Omega_{i}^{(j)}$ has a pole of order j|n|m+1j|n|-m+1 at $P_{a_{\nu}}$ with $\nu\neq i$ if $j|n|\geqslant m$ and of order $j|n|+1$ at $P_{a_{i}}.$ Note that, given that $n$ is coprime to $m$, the residues are nonzero only if $j$ is a multiple of $m$.
 Let us choose contours of integration in (8) in the form
-(38) $\gamma_{l-k}=\sum_{\nu=1}^{N}\frac{c^{\nu}_{l-k}}{2\pi\bf{\rm i}}\gamma_{a_{\nu}}$
+(38) $\gamma_{l-k}=\sum_{\nu=1}^{N}\frac{c^{\nu}_{l-k}}{2\pi{\mathrm{i}}}\gamma_{a_{\nu}}$
 where $c^{\nu}_{l-k}\in\mathbb{C}$ are arbitrary constants and where $\gamma_{a_{\nu}}$ is a small contour encircling the point $P_{a_{\nu}}$.
@@
 The nonzero residues from (39), for $(l-k)/m\in\mathbb{Z}$, are explicitly given by [8]
-(40) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=m\sum_{k_{1}+\ldots+k_{N}=(l-k)|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-a_{i})^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
+(40) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{i}^{(l-k)}(a)=m\hskip-28.45274pt\sum_{k_{1}+\ldots+k_{N}=(l-k)|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-a_{i})^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
 for $i=1,\ldots,N,\;i\neq\nu$ and
-(41) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{\nu}^{(l-k)}(a)=m\sum_{k_{1}+\ldots+\hat{k}_{\nu}+\ldots+k_{N}=(l-k)|n|/m}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
+(41) $\underset{P_{a_{\nu}}}{\rm res}\,\Omega_{\nu}^{(l-k)}(a)=m\hskip-28.45274pt\sum_{k_{1}+\ldots+\hat{k}_{\nu}+\ldots+k_{N}=(l-k)|n|/m}\prod_{h=1,h\neq\nu}^{N}\frac{{-(l-k)|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+(l-k)|n|/m}}$
 where the hat symbol over $k_{\nu}$ indicates that the summation index $k_{\nu}$ is absent in the sum in (41).
-Solutions obtained in Theorem 4 to the Fuchsian system (1) with polynomial coefficients (39) are rational functions of $z$ given by the next corollary.
+Solutions obtained in Theorem 4 to the Fuchsian system (1) with rational coefficients (39) are given by the next corollary.
 Corollary 10.
-Let $n<0,$ and $p,m,N>$ be integers such that ${\rm gcd}(n,m)=1$.
+Let $n<0,$ and $p,m,N>0$ be integers such that ${\rm gcd}(n,m)=1$.
 Let $\theta_{i}\in\mathbb{C}$ for $i=1,\dots,N$ and denote $\beta_{k}^{(i)}=\theta_{i}-\frac{(k-1)n}{m}$ for $k=1,\dots,p$.
-With $c_{j}\in\mathbb{C}$ as in (39), let $S_{j}=S_{j}(z)$ be the rational funtion
-(42) $S_{j}(z)=\begin{dcases*}\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\sum_{k_{1}+\ldots+k_{N}=j|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-z)^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-j|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+j|n|/m}},\quad\text{if }j/m\in\mathbb{Z}\\ \hskip 8.5359pt0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\text{if }j/m\notin\mathbb{Z}.\end{dcases*}$
+With $c_{j}^{\nu}\in\mathbb{C}$ as in (39), let $S_{j}=S_{j}(z)$ be the rational function
+(42) $S_{j}(z)=\begin{dcases*}\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\hskip-8.5359pt\sum_{k_{1}+\ldots+k_{N}=j|n|/m-1}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-z)^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-j|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+j|n|/m}},\quad\text{if }j/m\in\mathbb{Z}\\ \hskip 8.5359pt0,\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\quad\text{if }j/m\notin\mathbb{Z}.\end{dcases*}$
 Then the following upper triangular matrix $\Phi(z)$ solves the Fuchsian system (1) with coefficients $B^{(i)},\;i=1,\dots,N,$ being the upper triangular $p\times p$ matrices given by (39):
@@
 and
-(45) $M_{kl}=\begin{dcases*}\sum_{q\,\vdash\,(l-k)}\,\prod_{j=1}^{l-k}\frac{S_{j}^{\sigma_{q}(j)}}{\sigma_{q}(j)!}\quad&\text{when } $l>k$,\\ \quad 1\quad&\text{when } $l=k$,\\ \quad 0\quad&\text{when } $l<k$,\end{dcases*}$
+[…]
 where, as before, $\sigma_{q}(j)$ is the number of repetitions of the part $j$ in the partition $q$.
@@
 For $n<0$ and the compact Riemann surface $X_{a}$ corresponding to the superelliptic curve (4), the differentials $\Omega_{i}^{(j)}$ (6) have poles at $P_{a_{\nu}}$ for $\nu=1,\dots,N$ as shown at the beginning of Section 5.2.
-Similarly to the proof of Corollary 8, we need to prove that $\kappa_{j}$ from Definition 2 for the choice of the integration contour $\gamma_{j}$ as in (38) with $l-k=j$ coincides with $S_{j}$ (42).
+Similarly to the proof of Corollary 8, we need to prove that $\kappa_{j}$ from Definition 2 for the choice of the integration contour $\gamma_{j}$ as in (38) with lk=jl-k=j coincides with $S_{j}$ (42).
 We make an additional assumption that none of the contours $\gamma_{a_{\nu}}$ from (38) contains in its interior any point $P\in X_{a}$ for which $\zeta(P)=z$.
@@
 $\kappa_{j}=\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\underset{\xi_{a_{\nu}}=0}{\rm res}\;\frac{\Bigl{(}1+\frac{\xi_{a_{\nu}}^{m}}{a_{\nu}-z}\Bigr{)}^{-1}}{\xi_{a_{\nu}}^{j|n|-m+1}(a_{\nu}-z)}\prod_{h=1,h\neq\nu}^{N}\frac{d\xi_{a_{\nu}}}{(a_{\nu}-a_{h})^{j|n|/m}}\Bigl{(}1+\frac{\xi_{a_{\nu}}^{m}}{a_{\nu}-a_{h}}\Bigr{)}^{-j|n|/m}\\ =\frac{-m^{2}}{jn}\sum_{\nu=1}^{N}c^{\nu}_{j}\underset{\xi_{a_{\nu}}=0}{\rm res}\;\frac{d\xi_{a_{\nu}}}{\xi_{a_{\nu}}^{j|n|-m+1}}\sum_{r=0}^{\infty}\Bigl{[}\sum_{k_{1}+\ldots+k_{N}=r}\frac{(-1)^{k_{\nu}}}{(a_{\nu}-z)^{k_{\nu}+1}}\prod_{h=1,h\neq\nu}^{N}\frac{{-j|n|/m\choose k_{h}}}{(a_{\nu}-a_{h})^{k_{h}+j|n|/m}}\Bigr{]}\,\xi_{a_{\nu}}^{rm}.$
-Now, to obtain the residue, for each $\nu$, we take the term of the series in $\xi_{a_{\nu}}$ with $r$ such that $rm=j|n|-m$ if such an integer $r$ exists.
-If $j|n|-m$ is not an integer multiple of $m$, which is equivalent to $j$ not being an integer multiple of $m$ due to the choice of $n$ coprime with $m$, the residue vanishes.
+Now, to obtain the residue, for each $\nu$, we take the term of the series in $\xi_{a_{\nu}}$ with $r$ such that rm=j|n|mrm=j|n|-m if such an integer $r$ exists.
+If j|n|mj|n|-m is not an integer multiple of $m$, which is equivalent to $j$ not being an integer multiple of $m$ due to the choice of $n$ coprime with $m$, the residue vanishes.
 This gives us $S_{j}$ from the statement of the corollary.
@@
 Example 11.
-Let $p=3,\;m=N=2$, and $n=-1$.
+Let p=3,m=N=2p=3,\;m=N=2, and n=1n=-1.
 We thus consider the family of curves given by equation $w^{2}=(\zeta-a_{1})(\zeta-a_{2})$ where the associated Riemann surfaces are of genus zero.
-We define both integration contours to be a constant multiple of the small contour $\gamma_{a_{2}}$ around the point $P_{a_{2}}=(\zeta=a_{2},w=0)$ oriented counterclockwise on the surface: $\gamma_{1}=\gamma_{2}=\frac{c}{2\pi\bf{\rm i}}\gamma_{a_{2}}.$ Note that for every given $z\in\mathbb{C}$ the contour $\gamma_{a_{2}}$ is assumed not to encircle the points $P$ of the surface for which $\zeta(P)=z.$ For any $\theta_{1},\theta_{2},c\in\mathbb{C}$, Corollary 10 gives that the matrix
+We define both integration contours to be a constant multiple of the small contour $\gamma_{a_{2}}$ around the point $P_{a_{2}}=(\zeta=a_{2},w=0)$ oriented counterclockwise on the surface: $\gamma_{1}=\gamma_{2}=\frac{c}{2\pi{\mathrm{i}}}\gamma_{a_{2}}.$ Note that for every given $z\in\mathbb{C}$ the contour $\gamma_{a_{2}}$ is assumed not to encircle the points $P$ of the surface for which $\zeta(P)=z.$ For any $\theta_{1},\theta_{2},c\in\mathbb{C}$, Corollary 10 gives that the matrix
 $\Phi=\left(\begin{array}[]{ccc}(z-a_{1})^{\theta_{1}}(z-a_{2})^{\theta_{2}}&0&\frac{2c}{a_{1}-a_{2}}(z-a_{1})^{\theta_{1}+1}(z-a_{2})^{\theta_{2}}\\ 0&(z-a_{1})^{\theta_{1}+\frac{1}{2}}(z-a_{2})^{\theta_{2}+\frac{1}{2}}&0\\ 0&0&(z-a_{1})^{\theta_{1}+1}(z-a_{2})^{\theta_{2}+1}\end{array}\right)$
@@
 BP also acknowledges that this research was partially funded by the Fonds de recherche du Québec - section Nature et Technologies (FRQ).
+The authors thank Renat Gontsov for careful reading of the manuscript and pointing out some errors of notation and terminology, which greatly improved the presentation.
 VS gratefully acknowledges support from the Natural Sciences and Engineering Research Council of Canada through a Discovery grant and from the University of Sherbrooke.
+Data availability statement. The authors declare that there is no data associated with the manuscript.
+Conflict of interest statement. On behalf of all authors, the corresponding author states that there is no conflict of interest.